A Complete Bibliography of Philosophy of Science: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 November 2014 Version 1.01

Title word cross-reference

D [645]. ∈ [392].

-Thesis [645].

18-23 [312]. 1961 [312].

Abelson [477]. Abercrombie [175]. Abilities [210]. Abraham [88, 87]. Abstractive [324]. Abstracts [56, 9, 73, 74, 143, 126, 127, 164, 241, 255, 314, 313, 355, 405, 367, 366, 406, 462, 422, 461, 437, 463, 521, 537, 538, 504, 485, 505, 522, 539, 571, 553, 588, 554, 570, 589, 624, 649, 650, 623, 651]. Acceptance [632]. Account [634]. Ackermann [495]. Acquaintance [511, 391, 390]. Action [277, 16, 213]. Activity [516]. Adams [174]. Adaptation [428]. Adolf [608, 647]. against [147]. Ages [139]. Alan [31, 637]. Albert [420]. Albritton [620]. Alfred [179, 23, 333]. Alice [235].

Allan [330]. Alleged [250]. Allgemeinen [33]. Alternative [563]. Ambiguity [558]. Ambrose [235]. American [214]. among [44, 64, 395, 563]. Amplifying [527]. Analogies [352, 353]. Analogs [544]. Analogy [258, 259, 350, 351, 303]. Analyses [340]. Analysis [628, 14, 137, 307, 82, 237, 291]. Analytic [302, 195, 31]. Analytical [435, 535]. Analytische [70]. Anatomy [418, 175, 417]. Anaximander [176]. Andrew [333]. Angelelli [597]. Angell [421]. Animalivm [142]. Announcement [552, 609]. Anscombe [66]. Answers [80]. Anthropology [347, 261, 509, 50]. Antinomy [562]. Antiquity [48]. Anton [550]. Any [629]. Applications [184]. Applying [264]. Appraisal [39]. Approach [188, 254, 636]. Approaches [490]. Apuleian [567]. Apuleius [567]. Aren’t [232]. Argument [6, 394, 412, 411, 303, 45]. Arguments [301]. Aristotle [483, 198, 631, 646]. Arithmetic [419]. Arithmetization [102]. Arrow [612]. Arthur [271, 535, 71, 70]. Article [346]. Ascent [252]. Aspect [158]. Aspects [293, 413]. Assaying [97]. Association [360]. Astronomical [218]. Astronomy [198]. Asymmetric [170]. Authorities [201]. Authority [328]. Autocracy [177]. Axiologia [90]. Axiology [212]. Axiomatic [605, 162]. Axiomatischen [33]. Axiomatization [528, 395, 563]. Ayer [93].

B [421, 181, 20, 520, 218]. Babylon [240]. Back [10, 37, 57, 75, 94, 109, 128, 145, 165, 185, 203, 224, 242, 256, 273, 297, 315, 334, 348, 357, 368, 383, 408, 423, 449, 466, 486, 506, 523, 541, 555, 572, 590, 600, 610, 625, 638, 653]. Background [221]. Backward [508]. Bagby [156]. Baldwin [333]. Barberini [51]. Barnes [436]. Barnett [25]. Basic [619, 517]. Basis [209]. Be [452]. Beck [251]. Behavior [108, 5, 188, 15, 321, 55]. Behavioral [283]. Behaviorism [265]. Behavioristics [15]. Being [371]. Belief [123, 7, 515]. Bergmann [547]. Bernard [22]. Beth [419]. between [170]. Bibliography [220]. Biological [293, 428, 544, 492]. Biology [284, 345, 359, 115, 23, 617, 587]. Bird [138]. Birthday [202]. Black [279, 333, 475]. Board [360]. Bochenski [138]. Body [16]. Bonum [4]. Book [516, 550, 157, 200, 47, 284, 293, 235, 125, 238, 271, 535, 434, 598, 33, 141, 91, 330, 583, 66, 69, 221, 268, 67, 90, 216, 237, 480, 608, 24, 25, 183, 252, 253, 122, 421, 31, 196, 272, 295, 26, 46, 50, 119, 121, 156, 174, 199, 219, 220, 254, 267, 287, 291, 294, 331, 92, 106, 536, 88, 159, 161, 637, 517, 647, 478, 28, 53, 483, 27, 286, 290, 532, 107, 285, 327, 332, 93, 140, 382, 500, 551, 584, 582, 32, 179, 180, 236, 162, 178, 292, 68, 21, 123, 177]. Book [289, 23, 418, 534, 460, 329, 214, 585, 635, 30, 222, 158, 497, 498, 499, 622, 636, 634, 155, 181, 120, 435, 475, 476, 477, 482, 518, 566, 569, 20, 52, 197, 251, 269, 436, 549, 568, 19, 195, 502, 520, 548, 597, 419, 420, 34, 71, 213, 515, 533, 328, 270, 29, 163, 201, 503, 138, 160, 182, 70, 22, 35, 36, 48, 51, 54, 86, 139, 142, 176, 198, 215, 218, 240, 288, 417, 45, 87, 175, 333, 621, 479, 481, 501, 519, 567, 72, 89, 124, 184, 202, 217, 620, 49, 55]. Book [239, 587, 484, 586]. Boring [381]. Box [279]. Boyer [20]. Bradley [288]. Braithwaite [84]. Brenda [213]. Brian [517, 503]. Bridgman [21]. British [56, 314, 405, 462, 521, 537, 571, 624]. Brodbeck [448]. Bronfenbrenner [333]. Brown [50, 67, 478]. Buck [281]. Bunge [598, 107, 327, 85]. Burgers [516]. Burks [271]. Burrow [92].

C [238, 535, 480, 183, 435, 333, 481, 72]. Campbell [91]. Can [452]. Capek [236]. Carl [20, 270]. Carmody [218]. Carnap [157, 586, 402, 234, 637]. Carney [479]. Case [147]. Causal [107, 593, 2, 318]. Causality [345, 194, 85, 107]. Cause [453]. Caws [634]. Century [25]. Chains [278]. Challenges [234]. Chance [503]. Chandler [333]. Character [205, 509, 497]. Charles [271, 24, 199, 23, 549, 176, 49]. Chisholm [221]. Choice [510, 503]. Christensen [237]. Churchman [72, 238]. Cientifica [90]. City [568]. Civilizations [156]. Clagett [139, 139]. Claim [578]. Clarification [464, 640, 386]. Classic [183]. Classical [452, 48, 411]. Classics [284, 178, 201]. Clatterbaugh [391]. Claude [620]. Clause [529]. Clock [603]. Clocks [606]. Closed [123]. Cognitive [596, 121]. Cohen [22]. Coherence [580, 362]. Collected [271]. Collection [220, 51]. Colloquium [480]. Colloquy [312]. Commentary [265]. Comments [281, 192, 112, 299, 118, 171, 527, 307]. Commitments [338]. Committee [92]. Common [392]. Communication [529]. Companion [475]. Comparative [156]. Comparisons [322]. Complementarity [592, 607, 439]. Completeness [431, 534]. Complexity [247]. Computers [247]. Concept [200, 302, 18, 329, 213, 133]. Conception [159, 640, 641]. Conceptions [124]. Concepts [26, 517, 574, 595, 131]. Conceptual [516, 119, 120, 621]. Concerning [65, 250, 260, 578, 124]. Conclusive [643]. Condition [498]. Conditionals [317]. Conditions [431, 266]. Confirmation [227, 258, 578, 351, 560, 442]. Confirming [44, 64]. Conflict [496]. Connate [198]. Consistency [350]. Constants [410]. Construction [97]. Constructs [410]. Contemporary [236, 253, 87]. Context [602]. Contraction [642]. Contradiction [250]. Controversy [53]. Conventionality [489, 604, 605]. Coombs [457]. Coordinate [276]. Cornelis [288]. Corrected [432]. Correspondence [630]. Coser [267]. Cosmological [341, 440]. Cosmology [361, 180, 176]. Counterfactual [453, 317]. Counterfactuals [396]. Course [284, 201]. Crawshay [53]. Crawshay-Williams [53]. Criteria [114, 53]. Criterion [229]. Critical [433, 139]. Critique [365, 592, 607, 593, 124]. Culture [125, 156, 219].

D [216, 480, 520, 479, 474]. D-Thesis [474]. Daniel [160]. Danto [535]. D’Arcy [36, 36]. Darwin [25, 23, 24]. Data [457, 217]. David [71, 112]. Deception [293]. Deciding [81]. Decision [496, 238, 444, 443, 331, 498]. Decisions [44, 64]. Deducibility [370]. Deduction [431, 103]. Deductive [375, 579, 532, 266, 472]. Defense [329, 393]. Definition [132, 305]. Definitions [416, 1, 72, 397]. del [90]. Dematerialization [150, 149]. Democracy [177]. Derek [240]. Derived [628, 454]. Descent [18, 8]. Descriptive [493]. Design [47]. Determinants [214]. Determinism [454, 330]. Development [199]. Devereux [333]. Dialectica [9, 73, 143, 367, 422, 461, 538, 649]. Diamond [219]. Diesing [286]. Different [433]. Difficulties [525, 527]. Dimensional [628]. Dimensions [628, 354, 560, 291, 182]. Disabilities [210]. Disciplinarianism [343]. Disconfirmative [338]. Discourse [137]. Discovery [122]. Discrete [278]. Discussion [392, 433]. Disguised [325]. Dispositional [469]. Distance [213]. Distant [489]. Distinguishability [337]. Distributive [518]. Diversity [283]. Do [15]. Doctrine [456]. Documents [22]. Dorothy [26]. Dray [414]. Dreaming [69]. Drebbel [288]. Dual [205]. Dualism [179]. Duality [147]. Duhem [393, 644]. Duhemian [394, 6]. Duration [3]. Durkheim [220]. Dwight [23].

E. [381]. Eddington [124]. Edel [87]. edited [533]. Editorial [92]. Edward [382, 52, 333, 89, 202]. Effective [419]. Efficacy [318]. Egmont [237]. Einstein [249, 111, 460]. Elections [360]. Electromagnetism [39]. Elements [385, 500]. Elicit [15]. Ellis [517]. Emergence [456]. Emile [220]. Emmet [26]. Empedocle [198]. Empirical [100, 309]. Empiricist [62]. English [253]. Entities [342, 398]. Entrenchment [404]. Entropy [615, 380]. Epistemological [39]. Epistemology [345, 207]. Erkenntnislehre [70]. Ernest [233, 19]. Errata [108, 616]. Erratum [464]. Error [370, 293]. Erwin [484]. Essay [583, 216]. Essays [550, 293, 219, 220, 602, 585, 533, 202, 49]. Establishing [210]. Estimates [134]. Estimation [618]. Estructura [90]. etc [82]. Ethical [174]. Ethics [268, 270, 91, 178, 476, 477]. Eudoxan [198]. Eugene [533, 481]. Evaluating [217]. Events [172, 191]. Evert [419]. Evidence [172, 160]. Evolution [430, 387, 167, 252, 23]. Evolutionary [345, 59]. Examined [111]. Example [321]. Experience [480, 484, 516, 583, 31]. Experimental [451, 177, 370, 324, 217, 47]. Experimentalism [154]. Explanation [375, 432, 514, 135, 513, 170, 559, 231, 136, 266, 414, 359, 446, 455, 468, 413, 376, 278, 372, 472, 631, 646, 339, 478]. Explanation-Sketches [513]. Explanations [117]. Explanatory [319]. ’Explication [545, 545, 641]. Exposition [68, 124]. Extensionality [427]. Extensive [530].

Fabric [433]. Factors [444]. Facts [347]. Failing [388]. Fallacy [429]. Falling [580]. Falsification [643, 644]. Falsity [416, 642]. Feigl [65]. Felix [291, 158]. Feyerabend [458]. Feynman [497]. Field [32]. Fields [213]. Fieldwork [509]. Findlay [287]. Fire [215]. Fitzgerald [642]. Flour [48]. Fluents [228]. Foote [333]. Footnotes [617]. Force [627]. Forces [213]. Formal [235, 636, 419]. Formalists [637]. Formalization [400, 309]. Formulae [189]. Formulated [452]. Formulations [120, 119]. Fortbildung [34]. Foundation [92]. Foundations [15, 621, 605, 551]. Four [399]. Frame [289]. Frames [606]. Frameworks [168]. Francille [158]. Francis [218]. Frank [543, 17]. Frankena [476]. Fred [637]. Frederick [290]. Free [330]. Freed [122, 122]. Freedom [424, 291, 331]. Frege [520, 548, 597, 398, 502].

Freiberg [163]. Freud [329]. Front [11, 38, 58, 77, 95, 110, 129, 146, 166, 186, 204, 225, 243, 257, 274, 298, 316, 335, 349, 358, 369, 384, 409, 438, 450, 467, 487, 507, 524, 542, 556, 573, 591, 601, 611, 626, 639, 648]. Frontiers [52]. Function [121, 26]. Functional [359]. Functions [264]. Fundamental [395, 563]. Fundamentals [479]. Fundamentos [90]. Future [181].

G [66, 332, 32, 381, 620]. Galen [501]. Galton [79]. Gambling [153]. Game [491]. Games [633]. Gardner [199]. Garnett [91]. General [279, 390, 120]. Generalization [268]. Generalizations [353]. Geology [433]. Geometry [323]. Georg [46]. George [268, 482, 35, 586]. Gibb [337]. Gibbs [380]. Giorgio [285]. Giuseppe [263]. Glover [420]. Gödel [548, 104, 19]. Going [508]. Good [568]. Goodman [637, 616, 575, 404, 632]. Gottlob [597]. Goudge [252]. Governing [360]. Grain [48]. Grain-Mills [48]. Greek [54, 176, 215]. Grene [483]. Gross [27]. Growth [272]. Grue [616, 575]. Grünbaum [282, 593, 394, 640, 171, 308, 644, 641, 608, 647]. Grundlagen [33]. Gulliksen [184]. Gunpowder [215]. Gweneth [142].

H [88, 382, 158, 29, 176, 87, 89, 184]. Hacker [333]. Halberstadt [140]. Halmos [141]. Hanfmann [332]. Hans [267]. Hansen [124]. Hanson [150, 231]. Harald [331, 198]. Harris [288]. Hartman [90]. Harvey [142]. Hausman [637]. Haworth [568]. Hazel [436]. Heijenoort [548]. Hellensitic [35]. Helmut [294]. Hempel [135, 616, 575]. Henle [125]. Henry [636, 515, 45, 333]. Herman [550]. Hermeneias [567]. Hesse [213]. Hill [253]. Historical [550, 209, 551, 584]. Historical-Philosophical [209]. History [424, 14, 156, 219, 172, 471, 446, 71, 213, 35, 139, 215, 535, 29]. Hjalmar [435]. Hlavaty [158, 158]. Hogben [328]. Holdheim [267]. Honor [219]. Hook [182]. Horace [420]. Horizons [440]. Horowitz [239]. Hospers [365]. Houde [28]. Hughes [532]. Hugues [499]. Hull [29]. Human [88, 210, 321]. Humanist [289]. Humean [13]. Humphrey [288]. Hutchings [52]. Huxley [289]. Hypotheses [100, 643, 44, 64]. Hypothesis [642, 618].

Ideal [18, 8, 42]. Idealism [494]. Identity [512, 456]. Ignacio [597]. II [35, 607, 64, 246, 595, 646]. Image [292]. Images [183]. Impact [236]. Imperatives [63]. Important [552]. Imprecision [468]. Imre [635]. Incompatibilities [510]. Incomplete [398]. Indeterminacy [341]. Indirect [137]. Individuals [26]. Induction [41, 403, 441, 84, 105, 65, 130, 262, 336, 388, 317, 190]. Inductive [99, 259, 84, 499, 503]. Induktive [157]. Industry [214]. Inertia [244]. Inexact [385]. Inexactness [455]. Inference [235, 532, 160]. Infinite [361]. Influence [567]. Information [76, 144, 223, 296, 356, 407, 465, 540, 599, 652, 264, 211, 133]. Ingle [23]. Inquiry [74, 126, 164, 241, 255, 313, 355, 366, 406, 437, 463, 504, 553, 588, 650, 331, 53, 177, 418, 417]. Insurance [153]. Integration [254]. Intentions [287]. Interaction [447]. Interactionism [425]. Interdependence [14]. International [312]. Interpersonal [108, 55]. Interpretation [2]. Interrogatives [326, 82]. Intrinsic [640]. Introduction [66, 602, 140, 419, 96, 503, 621, 269, 536, 519]. Introductory [68]. Intuition [327]. Investigation [444]. Investigations [123]. Irving [239]. Isaac [22]. Israel [582, 418, 417]. Issues [238]. Ivor [68].

J [516, 293, 33, 287, 93, 23, 566, 533, 527, 138, 215, 218, 240]. Jack [584]. Jacob [121, 484]. Jaffe [159]. James [294, 19, 479]. Jan [500]. Jean [293, 436, 548]. Jean-Paul [436]. Jeremy [520]. Jerrold [566]. Johannes [124]. John [550, 622]. Johnson [175]. Johnstone [45]. Journal [56, 314, 405, 462, 521, 537, 571, 624]. Jr. [550]. Judgment [175]. Julian [289]. Julius [122, 158]. Justice [159, 518]. Justification [84]. Justifications [117].

Kahn [176]. Kant [562]. Kapp [254, 180]. Karl [122, 34]. Karlsson [46]. Katz [403, 566]. Kinds [114]. Kinship [116, 18]. Kleine [597]. Knowledge [88, 430, 615, 387, 202, 253, 254, 551, 584, 622, 239, 480]. Koch [119, 120]. Koopman [402]. Körner [583]. Kretzmann [519]. Krikorian [87]. Kritik [34]. Kuhn [473]. Kurt [220]. Kurtz [498]. Kyburg [636, 374, 515].

L [200, 295, 332, 29, 48, 288, 175, 333]. L. [527]. Lamprecht [421]. Lan [122]. Lancelot [328]. Landé [179]. Landsberger [333]. Language [18, 275, 8, 42, 211, 270, 200, 332, 329, 566, 125]. Languages [276]. Largely [150]. Latakos [635]. Latin [51]. Law [244, 290, 531, 497]. Lawlikeness [578, 559]. Lawrence [568]. Laws [114, 323, 445, 228, 324]. Lazerowitz [235]. Learning [120]. Leblanc [532, 499]. Leclerc [68]. Length [101]. Leonard [82]. Lerner [160]. Lesniewski [481]. Letters [92, 22]. Lewis [267]. Life [252, 293]. Lifwynn [92]. Limitations [400]. Linguistic [585]. Lippitt [177]. Literature [161]. Living [272]. Llewellyn [27]. Locali [142]. Loewenberg [121]. Logic [259, 43, 122, 557, 80, 445, 577, 399, 419, 515, 612, 411, 567, 202, 63, 421, 536, 28, 140, 500, 419, 503, 138, 479, 519, 235]. Logical [401, 614, 266, 7, 502, 15, 130, 481, 586, 93]. Logico [420]. Logico-Philosophical [420]. Logik [157]. Loose [350]. Lorentz [642, 189]. Lorentz-Formulae [189]. Louis [239]. Lucienne [158]. Luigi [55]. Lukasiewicz [500]. Luschei [481].

M [516, 330, 66, 221, 272, 50, 174, 199, 269, 195, 138, 175, 333, 310]. M-K [310]. Madden [382, 89]. Madison [329]. Makes [339]. Malcolm [69]. Man [183, 254, 294, 92, 292, 498, 181, 267, 292]. Mandelbaum [569]. Manier [430]. Manpower [214]. Marcson [214]. Marcucci [494]. Marcus [268]. Margenau [299]. Margolis [447]. Mario [598, 107, 327, 85]. Marjorie [483]. Mark [567]. Markov [278]. Markus [460]. Marshall [139]. Martin [195]. Marx [621]. Marxian [205]. Marxism [206, 346]. Marxist [424, 207].

Mary [213]. Maslow [88]. Massey [641]. Material [150]. Materialism [546, 59]. Mathematical [451, 263, 500, 340, 138]. Mathematics [635, 158, 20]. Matter [10, 37, 57, 75, 11, 38, 58, 77, 94, 109, 128, 145, 95, 110, 129, 146, 165, 185, 203, 224, 166, 186, 204, 225, 242, 256, 273, 297, 243, 257, 274, 298, 315, 334, 348, 357, 316, 335, 349, 358, 368, 383, 408, 369, 384, 409, 423, 449, 466, 438, 450, 467, 486, 506, 523, 541, 487, 507, 524, 542, 555, 572, 590, 600, 556, 573, 591, 601, 610, 625, 638, 653, 611, 626, 639, 648, 150, 149, 361]. Mattick [346]. Maurice [569]. Max [267, 475, 333]. Maximal [559, 558]. Maxwell [47]. May [161]. Mays [30]. MC [402]. McCrae [341]. Meaning [277, 152, 62]. Meanings [237]. Measurement [385, 581, 464, 628, 188, 12, 295, 187, 530, 102, 574, 595, 386, 517, 72]. Measurements [226, 246]. Measuring [363]. Mechanical [12]. Mechanics [452, 613, 139, 131, 411]. Mechanism [170]. Mechanisms [46, 308]. Medawar [181]. Medical [47]. Medicine [23]. Medieval [51]. Meeting [171]. Meiland [584]. Members [360]. Memory [543]. Menne [420]. Mental [546]. Messens [33]. Messick [184]. Metaethics [477]. Metamathematics [561]. Metaphysics [68, 216]. Method [585, 436, 50]. Methodological [488]. ’Methodology [576, 312, 163]. Methods [419, 71, 53]. Metric [640, 641]. Metrical [189]. Metricized [44, 64]. Meyerhoff [267]. Michael [533, 202]. Middle [139]. Milic [236]. Mill [190]. Mills [183, 48]. Milton [123]. Mind [16, 123, 182, 561, 287]. Minds [364]. Minkus [196]. Mistakes [151]. Mixture [198]. Modal [399]. Model [432, 264]. Models [352, 493]. Moderation [60]. Modern [174, 107, 140, 158, 608]. Modes [216]. Molecules [364]. Moore [533]. Moral [63]. Morality [290]. Moritz [48]. Morris [235, 549]. Morse [333]. Morton [551]. Motion [627]. Motv [142]. Mr. [346, 18, 430, 231, 65, 310]. Mukerjee [434, 197]. Müller [460]. Müller-Markus [460]. Multiplicative [395, 563]. Munn [330]. Murphy [199]. Murray [217]. Myhill [276]. Mystery [459]. Myth [20].

N [200, 50, 287, 251]. N. [150]. Nagel [19, 596, 233, 324]. Naive [141]. Names [392]. Naming [148]. Natural [627, 115, 23, 22]. Naturalism [174, 550]. Naturalist [36]. Nature [237, 247, 354, 261, 121, 123, 269, 567, 267, 620]. Necessary [210]. Necessity [378, 442, 169]. Needham [18]. Needleman [484]. Neglect [208]. Neglected [260]. Nelson [404]. Netherlanders [288]. Neumann [412]. Neuter [230]. Newman [513, 19]. Newton [22, 631, 646]. Newtonian [544]. Nicholas [536, 518, 501]. Niels [237]. Nomological [442]. Non [402, 411]. Non-Classical [411]. Non-Reducibility [402]. Noninertial [606]. Norman [69, 519]. Note [103, 82]. Notion [195]. Number [102]. Numerosity [102].

O [500, 180]. Oberlin [480]. Object [106]. Objective [594, 188]. Objectivism [447, 426, 425, 448]. Objectivity [441]. Obligation [63].

Observation [458, 229]. Observational [329]. Observations [208, 44, 64]. Officers [360]. Ofstad [331]. Olafson [290]. Oliver [292]. Ontological [130]. Ontology [391, 78, 390, 40, 547, 168, 234]. Open [123]. Operational [211]. Operations [419]. Operators [344]. Oppenheim [135, 291]. Optimal [238]. Order [227]. Organization [222, 211]. Organizational [214]. Origins [285, 176]. Orman [106]. Ostensive [1]. Otto [138]. Outlook [311].

P [550, 421, 196, 21, 533]. Panpsychism [59]. Pap [70]. Paper [281, 373, 112]. Papers [271, 22]. Paradigms [544]. Paradox [306, 380, 488, 633, 632, 337, 111]. Paradoxes [248, 613, 137, 442, 608]. Parameter [618]. Parity [170]. Parsons [333]. Part [592, 607, 64, 226, 246, 574, 595, 631, 646]. Partial [397]. Particular [320]. Particulars [511, 40, 564]. Partington [215]. Pasch [31]. Patrick [162]. Paul [125, 141, 219, 286, 498, 436]. Pauli [32]. Peano [263]. Peirce [271, 435, 153, 49]. Perception [108, 55, 569]. Perceptual [119, 199, 191]. Performance [343, 245]. Peri [567]. Periodicity [227, 259]. Permissible [464, 386]. Permission [245]. Person [108, 196, 55]. Personal [430, 526, 387, 525, 527, 202]. Personality [123, 434]. Perspectives [251]. Peter [329, 634, 181]. Petrullo [55]. Pfanzagl [33]. Phenomenalist [40]. Phenomenology [221]. Philburn [72]. Philip [156]. Philipp [543]. Philosopher [244]. Philosophers [54]. Philosophic [87]. Philosophical [238, 237, 252, 209, 236, 301, 585, 340, 647, 382, 420, 587]. Philosophy [424, 56, 127, 314, 405, 360, 462, 521, 537, 571, 624, 550, 61, 535, 434, 583, 325, 113, 287, 263, 154, 178, 381, 617, 635, 30, 634, 482, 566, 197, 251, 269, 17, 29, 22, 621, 124, 561, 49, 480, 121, 290, 586, 196, 636, 155, 569, 45, 239]. Phonons [562]. Physical [116, 347, 594, 323, 497, 337, 304, 124]. Physicalism [456]. Physics [78, 301, 206, 488, 275, 565, 179, 236, 213, 346, 86]. Physiological [119]. Pitcher [482]. Pitfalls [428]. Place [41, 267, 107]. Planning [619]. Plea [514]. Pneuma [198]. Podolsky [111]. Point [345]. Poland [312]. Polanyi [202]. Polarity [565]. Pomerans [293]. Popper [122]. Population [24]. Portrait [483]. Positive [254]. Positivism [93]. Positivist [249]. Postulate [412]. Postulates [152, 470]. Power [44, 64, 395, 563]. Powers [26]. Pragmatic [458, 159, 413]. Pragmatism [5, 435, 42]. Precis [138]. Precision [581]. Predicting [81]. Prediction [514, 431, 103, 170, 377, 231, 278, 472, 238]. Predictions [280, 579, 281]. Preference [230, 470, 612]. Preliminaries [302]. Present [3]. Presented [371, 202]. Presuppositions [317]. Price [240]. Primary [484]. Principle [341, 390, 107, 189, 620]. Principles [170, 23]. Prior [338, 372]. Prison [222]. Probabilistic [559, 411]. Probabilities [134, 372, 499]. Probability [528, 427, 452, 557, 526, 402, 443, 525, 527, 131, 362, 515]. Problem [216, 209, 531, 79, 7, 361, 173, 396]. Problem-Solving [209]. Problems [24, 647, 382, 13, 2, 278, 139, 587, 586, 87, 635]. Proceedings [480]. Processes [191, 120]. Professionalism [343]. Professor [494, 596, 281, 83, 192, 300, 171, 324, 308, 82]. Projecting [415]. Projection [412]. Prolegomena [156]. Proof [451, 454, 612, 19, 104]. Proofs [350]. Properties [469]. Proposal [260]. Proposed [43]. Propositional [399, 137]. Propositions [230]. Pseudo [586]. Pseudo-Problems [586]. Psi [173]. Psychoanalysis [127, 365, 446]. Psychological [13, 184]. Psychology [47, 488, 102, 382, 217, 119, 120]. Public [622]. Purpose [26, 245]. Putnam [192, 300, 104].

Qualitative [228]. Quantization [562]. Quantum [452, 12, 557, 342, 613, 147, 179, 301, 226, 246, 574, 595, 131, 411, 598]. Quasi [430]. Quasi-Reply [430]. Questions [232, 80, 304]. Quine [401, 106, 307, 644]. Qurra [218].

R [157, 141, 122, 150, 50, 19, 195, 215]. Radcliffe [50]. Radcliffe-Brown [50]. Radhakamal [434, 197]. Radical [265]. Radin [219]. Rainbow [20]. Ralph [177]. Ramsey [529, 410]. Ramsey-Sentence [529]. Randall [550]. Randomness [427]. Rational [414, 470, 515]. Ratoosh [72]. Ravens [459]. Raymond [159]. Raziel [477]. Re [111]. Re-Examined [111]. Reaction [447, 448]. Reading [378, 284]. Readings [290, 251, 201, 28]. Real [492]. Realism [221]. Realistic [526]. Reality [598]. Reason [514, 121, 286]. Reasoning [53, 4, 421]. Reducibility [402]. Reduction [132, 490]. Reference [18, 606]. Referring [216]. Reflections [121, 533]. Reflexive [280, 281, 377]. Reginald [180]. Reiche [198]. Reidentification [320]. Reiser [292]. Rejoinder [282, 194]. Related [22]. Relational [594]. Relations [512, 395, 563, 549]. Relationships [2]. Relativism [294]. Relativitatstheorie [34]. Relativity [439, 39, 603, 454, 605, 32]. Religion [161]. Remarks [276, 448, 275, 171, 362]. Renato [55]. Reply [391, 346, 430, 495, 83, 300, 448, 299, 374]. Repression [329]. Rescher [536, 518, 501]. Research [619, 23, 13, 217]. Resolution [306]. Response [150]. Ressentiment [267]. Results [360, 338]. Review [516, 550, 157, 200, 47, 284, 293, 235, 125, 238, 271, 535, 434, 598, 33, 141, 91, 330, 583, 66, 69, 221, 268, 67, 90, 216, 237, 480, 608, 24, 25, 183, 252, 253, 122, 421, 31, 196, 272, 295, 26, 46, 50, 119, 121, 156, 174, 199, 219, 220, 254, 287, 291, 294, 331, 92, 106, 536, 88, 159, 161, 637, 517, 647, 478, 28, 53, 483, 27, 286, 290, 532, 107, 285, 327, 332, 93, 140, 382, 500, 551, 584, 582, 32, 179, 180, 236, 162, 178, 292, 68, 21, 123, 177, 289]. Review [418, 534, 460, 329, 214, 585, 635, 30, 222, 158, 497, 498, 499, 622, 636, 634, 155, 181, 120, 435, 104, 475, 476, 477, 482, 518, 566, 569, 20, 52, 197, 251, 269, 436, 549, 568, 19, 195, 502, 520, 548, 597, 419, 420, 34, 71, 213, 515, 533, 328, 270, 29, 163, 201, 503, 138, 160, 182, 70, 22, 35, 36, 48, 51, 54, 86, 142, 176, 198, 215, 218, 240, 288, 417, 45, 87, 175, 333, 621, 479, 481, 501, 519, 567, 72, 89, 124, 184, 202, 217, 620, 49, 55, 239, 587]. Review [484, 586]. Reviews [267, 23, 139]. Revisited [136, 17, 472]. Revolution [43]. Revolutions [473]. Rex [54]. Richard [421, 534, 585, 497, 479]. Rigorous [454]. Risk [429]. Ritchie [71]. Robert [90, 478, 251, 502]. Robin [333]. Roderick [221]. Rods [363]. Roger [67, 281]. Rokeach [123]. Roland [28]. Role [167]. Rolf [586]. Rollins [480]. Rollo [161]. Ronald [177]. Rorty [585]. Rosen [111]. Rostand [293]. Rudolf [586]. Rules [630, 632]. Rupert [53]. Russell [23]. Ruth [36].

S [90, 216, 332, 460, 435, 420, 86, 184, 49]. S.A [25]. Safety [98]. Salmon [389, 373]. Sambursky [86]. Samuel [269]. Sanders [271]. Sanity [92]. Santillana [285]. Sapper [34]. Sarton [35]. Sartre [436]. Savage [527]. Saying [148]. Scaling [184]. Scheer [479]. Scheffler [582, 418, 417]. Scheler [267]. Schlegel [534]. Schlesinger [310]. Schoeck [294]. Scholar [36]. Scholar-Naturalist [36]. Schriften [597]. Science [56, 314, 405, 360, 462, 521, 537, 571, 624, 293, 238, 61, 41, 5, 247, 608, 325, 208, 119, 254, 209, 193, 212, 205, 233, 381, 636, 634, 120, 569, 17, 201, 173, 35, 139, 133, 239, 139, 583, 478, 107, 327, 534, 52, 197, 29, 35, 621, 124, 49, 582, 155, 328, 240]. Sciences [426, 14, 379, 312, 283, 565, 47, 71]. Scientific [375, 643, 61, 114, 97, 596, 444, 122, 113, 285, 509, 473, 278, 311, 62, 130, 4, 533, 163, 51, 621, 89, 217]. Scientist [229, 214]. Scriven [533]. Search [92, 436]. Second [562, 448]. Selected [284, 92, 201]. Selection [344, 23]. Semantic [428, 137]. Semanticists [344]. Semiorders [530]. Sense [569, 576]. Senses [484]. Sensory [119, 484]. Sentence [529]. Sept [312]. Sequences [427]. Seriousness [151]. Set [141, 162, 629]. Sets [510]. Seventieth [202]. Shandy [306]. Sharp [112]. Sherwood [519]. Shils [202]. Short [580]. Shwayder [216]. Sidman [217]. Sidney [182]. Sigmund [119, 120]. Significance [549, 309, 510]. Signification [549]. Signs [549]. Silverstein [51]. Simon [214]. Simple [399]. Simplicity [99, 260, 100, 97, 98, 187, 96, 307, 620]. Simulataneity [489]. Simultaneity [603, 604]. Simultaneous [528]. since [240]. Singer [268, 154]. Single [429]. Situation [577]. Skepticism [584]. Sketches [513, 376]. Skyrms [578, 503]. Slightly [526]. Slow [603, 606]. Smith [587]. Social [116, 347, 493, 426, 14, 261, 50, 254, 209, 478, 290, 205, 509, 222, 197, 333, 565, 46]. Societies [26]. Society [261, 254, 286, 290]. Sociological [183, 46, 27, 343]. Sociology [208, 239]. Solla [240]. Solley [199]. Solution [380]. Solutions [79]. Solving [209]. Some [302, 208, 444, 26, 531, 214, 2, 304, 307, 362]. Sound [562]. Sources [567]. Sowjetphilosophie [460]. Space [647, 564, 318]. Special [18, 120, 603, 454, 605]. Species [492]. Specificity [559, 558]. Specious [3]. Srinivas [50]. Stability [350]. Stanley [219]. State [278, 361]. Statements [232, 7]. States [226, 246, 191]. Statistical [558, 2, 372, 499]. Statistics [464, 386]. Status [596, 645, 372, 322]. Steady [361]. Steady-State [361]. Stegmüller [157]. Stephan [583]. Sterling [421]. Sternfeld [502]. Still [150]. Stimuli [15]. Stoics [86]. Strategy [619, 576]. Straus [484]. Strawson [564]. Strength [98]. Strict [580]. Structure [493, 193, 53, 233, 89, 586]. Studies [46, 222, 62, 420, 71]. Study [252, 26, 119, 156, 287, 294, 120, 549, 520, 419, 15, 435]. Styles [343]. Sub [342]. Sub-Quantum [342]. Subjective [528, 362]. Subjectivism [426].

Subjectivity [582]. Sullivan [567]. Summum [4]. Suppes [162]. Syllogism [501]. Symbolic [419]. Symbolism [161]. Symmetries [533]. Symmetry [231]. Symposium [620, 182, 27]. Synchronism [606]. Synonymy [401]. Synthese [485, 505, 522, 539, 554, 570, 589, 623, 651]. Synthetization [379]. System [402, 115]. Systematic [119, 634, 120]. Systems [12, 18, 123, 399, 8, 278, 337, 481, 272, 123].

T [252, 198]. Tactics [217]. Tagiuri [55]. Talcott [333]. Teaching [207]. Techniques [532]. Teleological [308]. Teleology [170]. Teleonomy [167]. Temporally [170]. Temporally-Asymmetric [170]. Terms [546, 397]. Testability [326]. Testing [488, 618]. Thabit [218]. Their [117]. Theodore [51]. Theodoric [163]. Theorems [402]. Theoretical [546, 329, 410, 574, 595, 397, 222]. Theorie [33]. Theories [352, 97, 596, 614, 253, 509, 168, 324, 333, 72]. Theory [227, 258, 385, 581, 464, 235, 598, 279, 594, 583, 458, 452, 252, 12, 557, 287, 471, 430, 603, 351, 404, 456, 593, 310, 387, 488, 305, 7, 278, 398, 454, 386, 62, 560, 525, 527, 457, 184, 629, 605, 141, 46, 27, 162, 502, 32]. Theory-Testing [488]. Thermodynamics [304]. These [602]. Thesis [645, 474]. Things [512, 121, 21, 67]. Thinking [183]. Thomas [253]. Thompson [36, 269, 36]. Thought [125, 621, 285, 89, 332]. Three [493, 614, 395, 563]. Thurstone [295]. Time [594, 508, 593, 647]. Tomas [49]. Touchstone [244]. Toulmin [631, 646]. Tractatus [66, 475]. Tradition [183]. Translations [220]. Transport [603, 606]. Trigant [92]. Tristram [306]. True [229, 629]. Truth [401, 416, 326, 230, 134, 228, 195]. Truth-Value [326, 134]. Truth-Valued [228]. Turn [585]. Two [637, 343, 613, 79, 288]. Typology [130].

Unconscious [302, 319]. Undecidable [304]. Understanding [550]. Unified [180]. Uniformity [620, 620]. Uniting [209]. Unity [179, 133]. Universals [216, 40]. Unpredictability [321]. Unprojectibles [415]. Urie [333]. Use [301]. Utility [528, 264]. Utilization [214].

Vakar [332]. Valor [90]. Value [287, 326, 134, 319]. Value-Theory [287]. Valued [228]. Values [238, 295, 88, 549, 287]. Variables [395, 563, 130]. Variety [258, 259]. Version [593]. versus [170]. View [345, 378, 174]. Viewpoint [211]. Vincent [49, 587]. Vindicate [388]. Vindicating [262]. Vindication [403, 105, 389, 65, 336, 484]. Volume [76, 144, 223, 296, 356, 407, 465, 540, 599, 652, 119, 120, 35]. vs [426, 618]. Vygotsky [332].

W [157, 271, 267, 294, 106, 584, 32, 21, 30, 419, 29, 45, 567]. Wahrscheinlichkeit [157]. Walker [520]. Wallace [23, 163]. Walsh [378]. Walter [533]. Warner [54]. Warsaw [312]. Wartofsky [621]. Way [21]. Weight [97]. Weingartner [118]. Welfare [322]. Wellman [270]. Wennerberg [435]. Wentworth [36]. Wesleyan [171]. West [238, 72].

Wheatley [83]. Whewell [378, 494, 190, 169]. White [177, 551]. Whitehead [30, 68]. Whitteridge [142]. Whyte [333]. Wiggins [294]. Wigner [299, 533]. Will [330]. William [254, 267, 140, 476, 163, 142, 333, 519]. Williams [53, 333]. Wilson [200, 637]. Witt [124]. Witt-Hansen [124]. Wittgenstein [66, 475, 482]. Wojtasiewicz [500]. Wolff [220]. Word [106]. Words [67]. Work [531]. Works [218]. World [24, 199, 484, 586, 174]. World-View [174]. Wright [183, 248]. Writings [51].

X [272].

Yervant [87].

Zarrow [272]. Zen [311]. Zeno [608]. Ziman [622].

References

Kotarbinska:1960:OD

[1] Janina Kotarbinska. On ostensive definitions. Philosophy of Science, 27 (1):1–22, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

     767X (electronic). URL http://www.jstor.org/stable/185302.

Nowak:1960:SPC

[2] Stefan Nowak. Some problems of causal interpretation of statistical relationships. Philosophy of Science, 27(1):23–38, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

     www.jstor.org/stable/185303.

Bergmann:1960:DSP

[3] Gustav Bergmann. Duration and the specious present. Philosophy of Science, 27(1):39–47, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

     185304.

Schlaretzki:1960:SRS

[4] W. E. Schlaretzki. Scientific reasoning and the summum bonum. Philosophy of Science, 27(1):48–57, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

     stable/185305.

Burns:1960:PSB

[5] Hobert W. Burns. Pragmatism and the science of behavior. Philosophy of Science, 27(1):58–74, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185306.

Grunbaum:1960:DA

[6] Adolf Grünbaum. The Duhemian argument. Philosophy of Science, 27(1): 75–87, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185307. See

comment [394].

Rescher:1960:PLT

[7] Nicholas Rescher. The problem of a logical theory of belief statements. Philosophy of Science, 27(1):88–95, January 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185308.

Needham:1960:DSI

[8] Rodney Needham. Descent systems and ideal language. Philosophy of Science, 27(1):96–101, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185309. See comments [18].

Anonymous:1960:ADa

[9] Anonymous. Abstracts from “dialectica”. Philosophy of Science, 27(1): 102–103, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185310.

Anonymous:1960:BMa

[10] Anonymous. Back matter. Philosophy of Science, 27(1):104–114, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185311.

Anonymous:1960:FMa

[11] Anonymous. Front matter. Philosophy of Science, 27(1):??, January 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185301.

Durand:1960:TMQ

[12] Loyal Durand, III. On the theory of measurement in quantum mechanical systems. Philosophy of Science, 27(2):115–133, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185887.

Naess:1960:PRH

[13] Siri Naess and Arne Naess. Psychological research and Humean problems. Philosophy of Science, 27(2):134–146, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185888.

Fisher:1960:AHI

[14] Franklin M. Fisher. On the analysis of history and the interdependence of the social sciences. Philosophy of Science, 27(2):147–158, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185889.

Rozeboom:1960:DSE

[15] William W. Rozeboom. Do stimuli elicit behavior? — A study in the logical foundations of behavioristics. Philosophy of Science, 27(2):159– 170, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185890.

Luce:1960:AMB

[16] David Randall Luce. The action of mind on body. Philosophy of Science, 27(2):171–182, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185891.

Rutherford:1960:FPS

[17] F. James Rutherford. Frank’s philosophy of science revisited. Philosophy of Science, 27(2):183–186, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185892.

Gellner:1960:CKS

[18] Ernest Gellner. The concept of kinship: With special reference to Mr. Needham’s “descent systems and ideal language”. Philosophy of Science, 27(2):187–204, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185893.

See [8].

Putnam:1960:BRB

[19] Hilary Putnam. Book review: Gödel’s Proof by Ernest Nagel; James R. Newman. Philosophy of Science, 27(2):205–207, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185894. See comments [104].

Ore:1960:BRB

[20] Oystein Ore. Book review: The Rainbow, from Myth to Mathematics by Carl B. Boyer. Philosophy of Science, 27(2):207–208, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185895.

Lucca:1960:BRB

[21] John De Lucca. Book review: The Way Things Are by P. W. Bridgman. Philosophy of Science, 27(2):208–209, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185896.

Stannard:1960:BRBa

[22] Jerry Stannard. Book review: Isaac Newton’s Papers & Letters on Natural Philosophy and Related Documents by I. Bernard Cohen. Philosophy of Science, 27(2):209–211, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185897.

Madden:1960:BRB

[23] Edward H. Madden. Book reviews: Evolution by Natural Selection by Charles Darwin; Alfred Russell Wallace; Principles of Research in Biology and Medicine by Dwight J. Ingle. Philosophy of Science, 27(2):211– 212, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185898.

Dickie:1960:BRBa

[24] George Dickie. Book review: The Problems of World Population by Charles Darwin. Philosophy of Science, 27(2):212–213, April 1960. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185899.

Dickie:1960:BRBb

[25] George Dickie. Book review: A Century of Darwin by S.A. Barnett. Philosophy of Science, 27(2):213–214, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185900.

Goldstein:1960:BRBa

[26] Leon J. Goldstein. Book review: Function, Purpose and Powers; Some Concepts in the Study of Individuals and Societies by Dorothy Emmet. Philosophy of Science, 27(2):214–217, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185901.

Haworth:1960:BRB

[27] Lawrence Haworth. Book review: Symposium on Sociological Theory by Llewellyn Gross. Philosophy of Science, 27(2):217, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185902.

Hardin:1960:BRBa

[28] Clyde Laurence Hardin. Book review: Readings in Logic by Roland Houde. Philosophy of Science, 27(2):217–218, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185903.

Shapere:1960:BRB

[29] Dudley Shapere. Book review: History and Philosophy of Science by L. W. H. Hull. Philosophy of Science, 27(2):218–220, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185904.

McClure:1960:BRB

[30] George McClure. Book review: The Philosophy of Whitehead by W. Mays. Philosophy of Science, 27(2):220–222, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185905.

Frankfurt:1960:BRB

[31] Harry G. Frankfurt. Book review: Experience and the Analytic by Alan Pasch. Philosophy of Science, 27(2):222–223, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185906.

Lenzen:1960:BRB

[32] V. F. Lenzen. Book review: Theory of Relativity by W. Pauli; G. Field. Philosophy of Science, 27(2):223–224, April 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185907.

Buchi:1960:BRB

[33] J. Richard Buchi. Book review: Die Axiomatischen Grundlagen Einer Allgemeinen Theorie des Messens by J. Pfanzagl. Philosophy of Science, 27(2):224–226, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185908.

Rosen:1960:BRBa

[34] Philip Rosen. Book review: Kritik und Fortbildung der Relativitatstheorie by Karl Sapper. Philosophy of Science, 27(2):226–227, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185909.

Stannard:1960:BRBb

[35] Jerry Stannard. Book review: A History of Science, Volume II, Hellensitic Science by George Sarton. Philosophy of Science, 27(2):227–228, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185910.

Stannard:1960:BRBc

[36] Jerry Stannard. Book review: D’Arcy Wentworth Thompson: The Scholar-Naturalist by Ruth D’Arcy Thompson. Philosophy of Science, 27(2):228–230, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185911.

Anonymous:1960:BMb

[37] Anonymous. Back matter. Philosophy of Science, 27(2):231–232, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185912.

Anonymous:1960:FMb

[38] Anonymous. Front matter. Philosophy of Science, 27(2):??, April 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185886.

Dingle:1960:REE

[39] Herbert Dingle. Relativity and electromagnetism: An epistemological appraisal. Philosophy of Science, 27(3):233–253, July 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185967.

Klemke:1960:UPP

[40] E. D. Klemke. Universals and particulars in a phenomenalist ontology. Philosophy of Science, 27(3):254–261, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185968.

Bunge:1960:PIS

[41] Mario Bunge. The place of induction in science. Philosophy of Science, 27(3):262–270, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185969.

Palmieri:1960:PIL

[42] L. E. Palmieri. Pragmatism and the ideal language. Philosophy of Science, 27(3):271–278, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185970.

Castaneda:1960:PRL

[43] Hector Neri Castaneda. On a proposed revolution in logic. Philosophy of Science, 27(3):279–292, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185971.

Finch:1960:CPOa

[44] Henry A. Finch. Confirming power of observations metricized for decisions among hypotheses. Philosophy of Science, 27(3):293–307, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185972.

Suits:1960:BRB

[45] Bernard Suits. Book review: Philosophy and Argument by Henry W. Johnstone. Philosophy of Science, 27(3):308–310, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185973.

Goldstein:1960:BRBb

[46] Leon J. Goldstein. Book review: Social Mechanisms, Studies in Sociological Theory by Georg Karlsson. Philosophy of Science, 27(3):310, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185974.

Barker:1960:BRB

[47] S. F. Barker. Book review: Experimental Design in Psychology and the Medical Sciences by A. E. Maxwell. Philosophy of Science, 27(3):310– 311, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185975.

Stannard:1960:BRBd

[48] Jerry Stannard. Book review: Grain-Mills and Flour in Classical Antiquity by L. A. Moritz. Philosophy of Science, 27(3):311–312, July 1960.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185976.

Wiener:1960:BRB

[49] Philip P. Wiener. Book review: Essays in the Philosophy of Science by Charles S. Peirce; Vincent Tomas. Philosophy of Science, 27(3):312–313, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185977.

Goldstein:1960:BRBc

[50] Leon J. Goldstein. Book review: Method in Social Anthropology by A. R. Radcliffe-Brown; M. N. Srinivas. Philosophy of Science, 27(3):313– 314, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185978.

Stannard:1960:BRBe

[51] Jerry Stannard. Book review: Medieval Latin Scientific Writings in the Barberini Collection by Theodore Silverstein. Philosophy of Science, 27 (3):314–315, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185979.

Palter:1960:BRB

[52] Robert M. Palter. Book review: Frontiers in Science by Edward Hutchings. Philosophy of Science, 27(3):316–319, July 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185980.

Hardin:1960:BRBb

[53] Clyde Laurence Hardin. Book review: Methods and Criteria of Reasoning: An Inquiry into the Structure of Controversy by Rupert Crawshay-Williams. Philosophy of Science, 27(3):319–320, July 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185981.

Stannard:1960:BRBf

[54] Jerry Stannard. Book review: The Greek Philosophers by Rex Warner. Philosophy of Science, 27(3):320–322, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185982.

Wispe:1960:BRB

[55] Lauren G. Wispe. Book review: Person Perception and Interpersonal Behavior by Renato Tagiuri; Luigi Petrullo. Philosophy of Science, 27(3):

322–324, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185983.

Anonymous:1960:ABJ

[56] Anonymous. Abstracts from “British Journal for the Philosophy of Science”. Philosophy of Science, 27(3):325–327, July 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185984.

Anonymous:1960:BMc

[57] Anonymous. Back matter. Philosophy of Science, 27(3):328, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185985.

Anonymous:1960:FMc

[58] Anonymous. Front matter. Philosophy of Science, 27(3):??, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185966.

Sellars:1960:PEM

[59] Roy Wood Sellars. Panpsychism or evolutionary materialism. Philosophy of Science, 27(4):329–350, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185235.

Edelheit:1960:M

[60] Rabbi Amos Edelheit. Moderation. Philosophy of Science, 27(4):350, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185236.

Benjamin:1960:PSS

[61] A. Cornelius Benjamin. Is the philosophy of science scientific? Philosophy of Science, 27(4):351–358, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185237.

Rozeboom:1960:SET

[62] William W. Rozeboom. Studies in the empiricist theory of scientific meaning. Philosophy of Science, 27(4):359–373, October 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185238.

Turnbull:1960:ILM

[63] Robert G. Turnbull. Imperatives, logic, and moral obligation. Philosophy of Science, 27(4):374–390, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185239.

Finch:1960:CPOb

[64] Henry A. Finch. Confirming power of observations metricized for decisions among hypotheses, Part II. Philosophy of Science, 27(4):391–404, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185240.

Kading:1960:CMF

[65] Daniel Kading. Concerning Mr. Feigl’s “vindication” of induction. Philosophy of Science, 27(4):405–407, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185241.

Carney:1960:BRBa

[66] James D. Carney. Book review: An Introduction to Wittgenstein’s Tractatus by G. E. M. Anscombe. Philosophy of Science, 27(4):408, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185242.

Carroll:1960:BRB

[67] John B. Carroll. Book review: Words and Things by Roger Brown. Philosophy of Science, 27(4):409–410, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185243.

Lowe:1960:BRB

[68] Victor Lowe. Book review: Whitehead’s Metaphysics: An Introductory Exposition by Ivor Leclerc. Philosophy of Science, 27(4):410–414, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185244.

Carney:1960:BRBb

[69] James D. Carney. Book review: Dreaming by Norman Malcolm. Philosophy of Science, 27(4):414–415, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185245.

Stadler:1960:BRB

[70] Ingrid Hess Stadler. Book review: Analytische Erkenntnislehre by Arthur Pap. Philosophy of Science, 27(4):415–418, October 1960. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185246.

Rosen:1960:BRBb

[71] Edward Rosen. Book review: Studies in the History and Methods of the Sciences by Arthur David Ritchie. Philosophy of Science, 27(4):418–420, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185247.

Thrall:1960:BRB

[72] R. M. Thrall. Book review: Measurement: Definitions and Theories by C. West Churchman; Philburn Ratoosh. Philosophy of Science, 27(4):420– 422, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185248.

Anonymous:1960:ADb

[73] Anonymous. Abstracts from “dialectica”. Philosophy of Science, 27(4): 423–424, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185249.

Anonymous:1960:AI

[74] Anonymous. Abstracts from “inquiry”. Philosophy of Science, 27(4):424– 427, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185250.

Anonymous:1960:BMd

[75] Anonymous. Back matter. Philosophy of Science, 27(4):428–434, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185251.

Anonymous:1960:VI

[76] Anonymous. Volume information. Philosophy of Science, 27(4):429–433, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185233.

Anonymous:1960:FMd

[77] Anonymous. Front matter. Philosophy of Science, 27(4):??, October 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185234.

Bergmann:1961:PO

[78] Gustav Bergmann. Physics and ontology. Philosophy of Science, 28(1):1– 14, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185460.

Naroll:1961:TSG

[79] Raoul Naroll. Two solutions to Galton’s problem. Philosophy of Science, 28(1):15–39, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185461.

Harrah:1961:LQA

[80] David Harrah. A logic of questions and answers. Philosophy of Science, 28(1):40–46, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185462.

Richfield:1961:DP

[81] Jerome Richfield and Irving M. Copi. Deciding and predicting. Philosophy of Science, 28(1):47–51, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185463.

Wheatley:1961:NPL

[82] J. M. O. Wheatley. Note on Professor Leonard’s analysis of interrogatives, etc. Philosophy of Science, 28(1):52–54, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185464.

Leonard:1961:RPW

[83] Henry S. Leonard. A reply to Professor Wheatley. Philosophy of Science, 28(1):55–64, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185465.

Coburn:1961:BIJ

[84] Robert C. Coburn. Braithwaite’s inductive justification of induction. Philosophy of Science, 28(1):65–71, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185466.

Schlegel:1961:MBC

[85] Richard Schlegel. Mario Bunge on causality. Philosophy of Science, 28 (1):72–82, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185467.

Stannard:1961:BRBa

[86] Jerry Stannard. Book review: Physics of the Stoics by S. Sambursky. Philosophy of Science, 28(1):83–84, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185468.

Suits:1961:BRB

[87] Bernard Suits. Book review: Contemporary Philosophic Problems by Yervant H. Krikorian; Abraham Edel. Philosophy of Science, 28(1):84–85, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185469.

Grabau:1961:BRB

[88] Richard F. Grabau. Book review: New Knowledge in Human Values by Abraham H. Maslow. Philosophy of Science, 28(1):85–86, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185470.

Tucker:1961:BRBa

[89] John Tucker. Book review: The Structure of Scientific Thought by Edward H. Madden. Philosophy of Science, 28(1):86–89, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185471.

Castaneda:1961:BRB

[90] Hector Castaneda. Book review: La Estructura del Valor: Fundamentos de la Axiologia Cientifica by Robert S. Hartman. Philosophy of Science, 28(1):89–93, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185472.

Burgener:1961:BRB

[91] R. J. C. Burgener. Book review: Ethics by A. Campbell Garnett. Philosophy of Science, 28(1):93–94, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185473.

Golightly:1961:BRB

[92] Cornelius L. Golightly. Book review: A Search for Man’s Sanity: The Selected Letters of Trigant Burrow by Editorial Committee of the Lifwynn Foundation. Philosophy of Science, 28(1):94–95, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185474.

Johnstone:1961:BRBa

[93] Henry W. Johnstone, Jr. Book review: Logical Positivism by A. J. Ayer. Philosophy of Science, 28(1):95–96, January 1961. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185475.

Anonymous:1961:BMa

[94] Anonymous. Back matter. Philosophy of Science, 28(1):97–108, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185476.

Anonymous:1961:FMa

[95] Anonymous. Front matter. Philosophy of Science, 28(1):??, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185459.

Rudner:1961:IS

[96] Richard Rudner. An introduction to simplicity. Philosophy of Science, 28(2):109–119, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185803.

Bunge:1961:WSC

[97] Mario Bunge. The weight of simplicity in the construction and assaying of scientific theories. Philosophy of Science, 28(2):120–149, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185804.

Goodman:1961:SSS

[98] Nelson Goodman. Safety, strength, simplicity. Philosophy of Science, 28(2):150–151, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185805.

Ackermann:1961:IS

[99] Robert Ackermann. Inductive simplicity. Philosophy of Science, 28(2): 152–161, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185806.

Barker:1961:SEH

[100] S. F. Barker. On simplicity in empirical hypotheses. Philosophy of Science, 28(2):162–171, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185807.

Henmueller:1961:WL

[101] Frank Henmueller and Karl Menger. What is length? Philosophy of Science, 28(2):172–177, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185808.

Nelson:1961:NNA

[102] Thomas M. Nelson and S. Howard Bartley. Numerosity, number, arithmetization, measurement and psychology. Philosophy of Science, 28(2): 178–203, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185809.

Canfield:1961:NPD

[103] John Canfield and Keith Lehrer. A note on prediction and deduction. Philosophy of Science, 28(2):204–208, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185810.

Nagel:1961:PRB

[104] Ernest Nagel and James R. Newman. Putnam’s review of Gödel’s Proof. Philosophy of Science, 28(2):209–211, April 1961. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185811. See [19].

Feigl:1961:VI

[105] Herbert Feigl. On the vindication of induction. Philosophy of Science, 28(2):212–216, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185812.

Goodstein:1961:BRB

[106] R. L. Goodstein. Book review: Word and Object by W. Van Orman Quine. Philosophy of Science, 28(2):217, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185813.

Hinshaw:1961:BRB

[107] Virgil Hinshaw, Jr. Book review: Causality: The Place of the Causal Principle in Modern Science by Mario Bunge. Philosophy of Science, 28(2):218–222, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185814.

Anonymous:1961:EPP

[108] Anonymous. Errata: Person perception and interpersonal behavior. Philosophy of Science, 28(2):222, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185815.

Anonymous:1961:BMb

[109] Anonymous. Back matter. Philosophy of Science, 28(2):223, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185816.

Anonymous:1961:FMb

[110] Anonymous. Front matter. Philosophy of Science, 28(2):??, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185802.

Sharp:1961:EPR

[111] David H. Sharp. The Einstein–Podolsky–Rosen Paradox re-examined. Philosophy of Science, 28(3):225–233, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185108.

Putnam:1961:CPD

[112] Hilary Putnam. Comments on the paper of David sharp. Philosophy of Science, 28(3):234–237, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185109.

Feibleman:1961:SP

[113] James K. Feibleman. The scientific philosophy. Philosophy of Science, 28(3):238–259, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185110.

Bunge:1961:KCS

[114] Mario Bunge. Kinds and criteria of scientific laws. Philosophy of Science, 28(3):260–281, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185111.

Lorch:1961:NSB

[115] J. Lorch. The natural system in biology. Philosophy of Science, 28(3):282– 295, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185112.

Barnes:1961:PSK

[116] J. A. Barnes. Physical and social kinship. Philosophy of Science, 28(3): 296–299, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185113.

Weingartner:1961:ETJ

[117] Rudolph H. Weingartner. Explanations and their justifications. Philosophy of Science, 28(3):300–305, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185114.

Scriven:1961:CW

[118] Michael Scriven. Comments on Weingartner. Philosophy of Science, 28 (3):306, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185115.

Goldstein:1961:BRBa

[119] Alvin G. Goldstein. Book review: Psychology: A Study of a Science. Study I, Conceptual and Systematic. Volume 1, Sensory, Perceptual and Physiological Formulations by Sigmund Koch. Philosophy of Science, 28 (3):307–311, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185116.

Mowrer:1961:BRB

[120] O. Hobart Mowrer. Book review: Psychology: A Study of a Science. Study I, Conceptual and Systematic. Volume 2, General Systematic Formulations, Learning, and Special Processes by Sigmund Koch. Philosophy of Science, 28(3):312–317, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185117.

Goldstein:1961:BRBb

[121] Leon J. Goldstein. Book review: Reason and the Nature of Things: Reflections on the Cognitive Function of Philosophy by Jacob Loewenberg. Philosophy of Science, 28(3):317–319, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185118.

Fain:1961:BRB

[122] Haskell Fain. Book review: The Logic of Scientific Discovery by Karl R. Popper; Julius Freed; Lan Freed. Philosophy of Science, 28(3):319– 324, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185119.

MacKinnon:1961:BRB

[123] William J. MacKinnon. Book review: The Open and Closed Mind: Investigations into the Nature of Belief Systems and Personality Systems by Milton Rokeach. Philosophy of Science, 28(3):324–327, July 1961. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185120.

Tucker:1961:BRBb

[124] John Tucker. Book review: Exposition and Critique of the Conceptions of Eddington concerning the Philosophy of Physical Science by Johannes Witt-Hansen. Philosophy of Science, 28(3):327–328, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185121.

Bedau:1961:BRB

[125] Hugo A. Bedau. Book review: Language, Thought & Culture by Paul Henle. Philosophy of Science, 28(3):328–330, July 1961. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185122.

Anonymous:1961:AI

[126] Anonymous. Abstracts from “inquiry”. Philosophy of Science, 28(3):331– 333, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185123.

Anonymous:1961:APP

[127] Anonymous. Abstracts from “philosophy and psychoanalysis”. Philosophy of Science, 28(3):333–334, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185124.

Anonymous:1961:BMc

[128] Anonymous. Back matter. Philosophy of Science, 28(3):335, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185125.

Anonymous:1961:FMc

[129] Anonymous. Front matter. Philosophy of Science, 28(3):??, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185107.

Rozeboom:1961:OIL

[130] William W. Rozeboom. Ontological induction and the logical typology of scientific variables. Philosophy of Science, 28(4):337–377, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185479.

Suppes:1961:PCQ

[131] Patrick Suppes. Probability concepts in quantum mechanics. Philosophy of Science, 28(4):378–389, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185480.

Madden:1961:DR

[132] Edward H. Madden. Definition and reduction. Philosophy of Science, 28 (4):390–405, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185481.

Wilkinson:1961:CIU

[133] John Wilkinson. The concept of information and the unity of science. Philosophy of Science, 28(4):406–413, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185482.

Leblanc:1961:PTV

[134] Hugues Leblanc. Probabilities as truth-value estimates. Philosophy of Science, 28(4):414–417, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185483.

Eberle:1961:HOE

[135] Rolf Eberle, David Kaplan, and Richard Montague. Hempel and Oppenheim on explanation. Philosophy of Science, 28(4):418–428, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185484.

Kaplan:1961:ER

[136] David Kaplan. Explanation revisited. Philosophy of Science, 28(4):429– 436, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185485.

Rescher:1961:SPP

[137] Nicholas Rescher. Semantic paradoxes and the propositional analysis of indirect discourse. Philosophy of Science, 28(4):437–440, October 1961.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185486.

Simons:1961:BRB

[138] Leo Simons. Book review: A Precis of Mathematical Logic by J. M. Bochenski; Otto Bird. Philosophy of Science, 28(4):441–442, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185487.

Stannard:1961:BRBb

[139] Jerry Stannard. Book reviews: The Science of Mechanics in the Middle Ages by Marshall Clagett; Critical Problems in the History of Science by Marshall Clagett. Philosophy of Science, 28(4):442–444, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185488.

Johnstone:1961:BRBb

[140] Henry W. Johnstone, Jr. Book review: An Introduction to Modern Logic by William Halberstadt. Philosophy of Science, 28(4):444, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185489.

Buchi:1961:BRB

[141] J. Richard Buchi. Book review: Naive Set Theory by Paul R. Halmos. Philosophy of Science, 28(4):445, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185490.

Stannard:1961:BRBc

[142] Jerry Stannard. Book review: De Motv Locali Animalivm by William Harvey; Gweneth Whitteridge. Philosophy of Science, 28(4):445–446, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185491.

Anonymous:1961:AD

[143] Anonymous. Abstracts from dialectica. Philosophy of Science, 28(4):447– 448, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185492.

Anonymous:1961:VI

[144] Anonymous. Volume information. Philosophy of Science, 28(4):449–453, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185477.

Anonymous:1961:BMd

[145] Anonymous. Back matter. Philosophy of Science, 28(4):454, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185493.

Anonymous:1961:FMd

[146] Anonymous. Front matter. Philosophy of Science, 28(4):??, October 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185478.

Lande:1962:CAQ

[147] Alfred Landé. The case against quantum duality. Philosophy of Science, 29(1):1–6, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185172.

Sellars:1962:NS

[148] Wilfrid Sellars. Naming and saying. Philosophy of Science, 29(1):7–26, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185173.

Hanson:1962:DM

[149] Norwood Russell Hanson. The dematerialization of matter. Philosophy of Science, 29(1):27–38, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185174. See response [150].

Feigl:1962:MSL

[150] Herbert Feigl. Matter still largely material. (a response to N. R. Hanson’s “The Dematerialization of Matter”). Philosophy of Science, 29(1):39–46, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185175. See [149].

Levi:1962:SM

[151] Isaac Levi. On the seriousness of mistakes. Philosophy of Science, 29 (1):47–65, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185176.

Madden:1962:PM

[152] Edward H. Madden and Murray J. Kiteley. Postulates and meaning. Philosophy of Science, 29(1):66–78, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185177.

Ullian:1962:PGI

[153] Joseph S. Ullian. Peirce, gambling, and insurance. Philosophy of Science, 29(1):79–80, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185178.

Krikorian:1962:SPE

[154] Y. H. Krikorian. Singer’s philosophy of experimentalism. Philosophy of Science, 29(1):81–91, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185179.

Moore:1962:BRBa

[155] Edward C. Moore. Book review: Philosophy of Science. Philosophy of Science, 29(1):92–93, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185180.

Goldstein:1962:BRBa

[156] Leon J. Goldstein. Book review: Culture and History, Prolegomena to the Comparative Study of Civilizations by Philip Bagby. Philosophy of Science, 29(1):93–94, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185181.

Bar-Hillel:1962:BRBa

[157] Y. Bar-Hillel. Book review: Induktive Logik und Wahrscheinlichkeit by R. Carnap; W. Stegmüller. Philosophy of Science, 29(1):94–95, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185182.

Menger:1962:BRB

[158] Karl Menger. Book review: The Modern Aspect of Mathematics by Lucienne Félix; Julius H. Hlavaty; Francille H. Hlavaty. Philosophy of Science, 29(1):95–96, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185183.

Grabau:1962:BRBa

[159] Richard F. Grabau. Book review: The Pragmatic Conception of Justice by Raymond Jaffe. Philosophy of Science, 29(1):96–98, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185184.

Simons:1962:BRB

[160] Leo Simons. Book review: Evidence and Inference by Daniel Lerner. Philosophy of Science, 29(1):98, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185185.

Grabau:1962:BRBb

[161] Richard F. Grabau. Book review: Symbolism in Religion and Literature by Rollo May. Philosophy of Science, 29(1):98–99, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185186.

Levy:1962:BRB

[162] Azriel Levy. Book review: Axiomatic Set Theory by Patrick Suppes. Philosophy of Science, 29(1):99–101, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185187.

Shapere:1962:BRBa

[163] Dudley Shapere. Book review: The Scientific Methodology of Theodoric of Freiberg by William A. Wallace. Philosophy of Science, 29(1):101–102, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185188.

Anonymous:1962:AI

[164] Anonymous. Abstracts from inquiry. Philosophy of Science, 29(1):103– 104, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185189.

Anonymous:1962:BMa

[165] Anonymous. Back matter. Philosophy of Science, 29(1):105–116, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185190.

Anonymous:1962:FMa

[166] Anonymous. Front matter. Philosophy of Science, 29(1):??, January 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185171.

deLaguna:1962:RTE

[167] Grace A. de Laguna. The role of teleonomy in evolution. Philosophy of Science, 29(2):117–131, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186539.

Maxwell:1962:TFO

[168] Grover Maxwell. Theories, frameworks, and ontology. Philosophy of Science, 29(2):132–138, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186540.

Walsh:1962:WN

[169] Harold T. Walsh. Whewell on necessity. Philosophy of Science, 29(2):139– 145, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186541.

Grunbaum:1962:TAP

[170] Adolf Grünbaum. Temporally-asymmetric principles, parity between explanation and prediction, and mechanism versus teleology. Philosophy of Science, 29(2):146–170, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186542.

Scriven:1962:CPG

[171] Michael Scriven. Comments on Professor Grünbaum’s remarks at the Wesleyan Meeting. Philosophy of Science, 29(2):171–174, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186543.

Goldstein:1962:EEH

[172] Leon J. Goldstein. Evidence and events in history. Philosophy of Science, 29(2):175–194, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186544.

Shewmaker:1962:SPP

[173] Kenneth L. Shewmaker and Carlton W. Berenda. Science and the problem of psi. Philosophy of Science, 29(2):195–203, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186545.

Goldstein:1962:BRBb

[174] Leon J. Goldstein. Book review: Ethical Naturalism and the Modern World-View by E. M. Adams. Philosophy of Science, 29(2):204–206, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186546.

Suits:1962:BRB

[175] Bernard Suits. Book review: The Anatomy of Judgment by M. L. Johnson Abercrombie. Philosophy of Science, 29(2):206–207, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186547.

Stannard:1962:BRBa

[176] Jerry Stannard. Book review: Anaximander and the Origins of Greek Cosmology by Charles H. Kahn. Philosophy of Science, 29(2):207–209, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186548.

MacKinnon:1962:BRB

[177] William J. MacKinnon. Book review: Autocracy and Democracy: An Experimental Inquiry by Ralph White; Ronald Lippitt. Philosophy of Science, 29(2):209–212, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186549.

Linden:1962:BRB

[178] George W. Linden. Book review: Classics in Philosophy and Ethics. Philosophy of Science, 29(2):212–213, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186550.

Lenzen:1962:BRBa

[179] V. F. Lenzen. Book review: From Dualism to Unity in Quantum Physics by Alfred Landé. Philosophy of Science, 29(2):213–216, April 1962. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186551.

Lenzen:1962:BRBb

[180] V. F. Lenzen. Book review: Towards a Unified Cosmology by Reginald O. Kapp. Philosophy of Science, 29(2):216–217, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186552.

Moore:1962:BRBb

[181] Edward C. Moore. Book review: The Future of Man by Peter B. Medawar. Philosophy of Science, 29(2):217–218, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186553.

Sosensky:1962:BRB

[182] Irving Sosensky. Book review: Dimensions of Mind, a Symposium by Sidney Hook. Philosophy of Science, 29(2):218–220, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186554.

Dickie:1962:BRB

[183] George Dickie. Book review: Images of Man: The Classic Tradition in Sociological Thinking by C. Wright Mills. Philosophy of Science, 29(2): 220–221, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186555.

Tucker:1962:BRBa

[184] John Tucker. Book review: Psychological Scaling: Theory and Applications by H. Gulliksen; S. Messick. Philosophy of Science, 29(2):221, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186556.

Anonymous:1962:BMb

[185] Anonymous. Back matter. Philosophy of Science, 29(2):222–223, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186557.

Anonymous:1962:FMb

[186] Anonymous. Front matter. Philosophy of Science, 29(2):??, April 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186538.

Hillman:1962:MS

[187] Donald J. Hillman. The measurement of simplicity. Philosophy of Science, 29(3):225–252, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186043.

Delgado:1962:OAM

[188] R. Rodriguez Delgado and J. M. R. Delgado. An objective approach to measurement of behavior. Philosophy of Science, 29(3):253–268, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186044.

Tornebohm:1962:LFM

[189] Håkan Törnebohm. The Lorentz-formulae and the metrical principle. Philosophy of Science, 29(3):269–278, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186045.

Walsh:1962:WMI

[190] Harold T. Walsh. Whewell and Mill on induction. Philosophy of Science, 29(3):279–284, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186046.

Myers:1962:PES

[191] Charles Mason Myers. Perceptual events, states, and processes. Philosophy of Science, 29(3):285–291, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186047.

Margenau:1962:CPP

[192] H. Margenau and E. P. Wigner. Comments on Professor Putnam’s comments. Philosophy of Science, 29(3):292–293, July 1962. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186048.

Grunbaum:1962:SS

[193] Adolf Grünbaum. The structure of science. Philosophy of Science, 29(3): 294–305, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186049.

Bunge:1962:CR

[194] Mario Bunge. Causality: A rejoinder. Philosophy of Science, 29(3):306– 317, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186050.

Putnam:1962:BRB

[195] Hilary Putnam. Book review: The Notion of Analytic Truth by R. M. Martin. Philosophy of Science, 29(3):318–320, July 1962. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186051.

Ginet:1962:BRB

[196] Carl Ginet. Book review: Philosophy of the Person by P. A. Minkus. Philosophy of Science, 29(3):320–322, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186052.

Parsons:1962:BRB

[197] Howard L. Parsons. Book review: The Philosophy of Social Science by Radhakamal Mukerjee. Philosophy of Science, 29(3):322–323, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186053.

Stannard:1962:BRBb

[198] Jerry Stannard. Book review: Empedocles’ Mixture, Eudoxan Astronomy and Aristotle’s Connate Pneuma by Harald A. T. Reiche. Philosophy of Science, 29(3):323–324, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186054.

Goldstein:1962:BRBc

[199] Alvin G. Goldstein. Book review: Development of the Perceptual World by Charles M. Solley; Gardner Murphy. Philosophy of Science, 29(3):324– 326, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186055.

Bar-Hillel:1962:BRBb

[200] Y. Bar-Hillel. Book review: The Concept of Language by N. L. Wilson. Philosophy of Science, 29(3):326, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186056.

Shapere:1962:BRBb

[201] Dudley Shapere. Book review: Classics in Science: A Course of Selected Readings by Authorities. Philosophy of Science, 29(3):326–328, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186057.

Tucker:1962:BRBb

[202] John Tucker. Book review: The Logic of Personal Knowledge, Essays Presented to Michael Polanyi on His Seventieth Birthday by Edward Shils. Philosophy of Science, 29(3):328–329, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186058.

Anonymous:1962:BMc

[203] Anonymous. Back matter. Philosophy of Science, 29(3):330–331, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186059.

Anonymous:1962:FMc

[204] Anonymous. Front matter. Philosophy of Science, 29(3):??, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186042.

Hodges:1962:DCM

[205] Donald Clark Hodges. The dual character of Marxian social science. Philosophy of Science, 29(4):333–349, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186293.

Mattick:1962:MNP

[206] Paul Mattick. Marxism and the new physics. Philosophy of Science, 29 (4):350–364, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186294.

See reply [346].

Dunham:1962:TME

[207] Barrows Dunham. On teaching Marxist epistemology. Philosophy of Science, 29(4):365–368, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186295.

Ehrlich:1962:SON

[208] Howard J. Ehrlich. Some observations on the neglect of the sociology of science. Philosophy of Science, 29(4):369–376, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186296.

Goodwin:1962:HPB

[209] Leonard Goodwin. The historical-philosophical basis for uniting social science with social problem-solving. Philosophy of Science, 29(4):377– 392, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186297.

McGill:1962:ENH

[210] V. J. McGill. On establishing necessary human abilities and disabilities. Philosophy of Science, 29(4):393–405, October 1962. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186298.

Rothstein:1962:IOL

[211] Jerome Rothstein. Information and organization as the language of the operational viewpoint. Philosophy of Science, 29(4):406–411, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186299.

Hartman:1962:AS

[212] Robert S. Hartman. Axiology as a science. Philosophy of Science, 29 (4):412–433, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186300.

Rosen:1962:BRB

[213] Edward Rosen. Book review: Forces and Fields: The Concept of Action at a Distance in the History of Physics by Mary Brenda Hesse. Philosophy of Science, 29(4):434–435, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186301.

Mayer:1962:BRB

[214] Kurt B. Mayer. Book review: The Scientist in American Industry: Some Organizational Determinants in Manpower Utilization by Simon Marcson. Philosophy of Science, 29(4):435–436, October 1962. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186302.

Stannard:1962:BRBc

[215] Jerry Stannard. Book review: A History of Greek Fire and Gunpowder by J. R. Partington. Philosophy of Science, 29(4):436–438, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186303.

Caton:1962:BRB

[216] Charles E. Caton. Book review: Modes of Referring and the Problem of Universals, an Essay in Metaphysics by D. S. Shwayder. Philosophy of Science, 29(4):438–439, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186304.

Wasson:1962:BRB

[217] Chester R. Wasson. Book review: Tactics of Scientific Research: Evaluating Experimental Data in Psychology by Murray Sidman. Philosophy of Science, 29(4):439–441, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186305.

Stannard:1962:BRBd

[218] Jerry Stannard. Book review: The Astronomical Works of Thabit b. Qurra by Francis J. Carmody. Philosophy of Science, 29(4):441, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186306.

Goldstein:1962:BRBd

[219] Leon F. Goldstein. Book review: Culture in History: Essays in Honor of Paul Radin by Stanley Diamond. Philosophy of Science, 29(4):442–443, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186307.

Goldstein:1962:BRBe

[220] Leon J. Goldstein. Book review: Emile Durkheim, 1858–1917; A Collection of Essays with Translations and Bibliography by Kurt Wolff. Philosophy of Science, 29(4):443–444, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186308.

Carney:1962:BRB

[221] James D. Carney. Book review: Realism and the Background of Phenomenology by Roderick M. Chisholm. Philosophy of Science, 29(4):444– 445, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186309.

McClure:1962:BRB

[222] George McClure. Book review: Theoretical Studies in Social Organization of the Prison. Philosophy of Science, 29(4):446, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186310.

Anonymous:1962:VI

[223] Anonymous. Volume information. Philosophy of Science, 29(4):447–451, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186291.

Anonymous:1962:BMd

[224] Anonymous. Back matter. Philosophy of Science, 29(4):452–453, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186311.

Anonymous:1962:FMd

[225] Anonymous. Front matter. Philosophy of Science, 29(4):??, October 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186292.

Margenau:1963:MQSa

[226] Henry Margenau. Measurements and quantum states: Part I. Philosophy of Science, 30(1):1–16, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186615.

Achinstein:1963:CTO

[227] Peter Achinstein. Confirmation theory, order, and periodicity. Philosophy of Science, 30(1):17–35, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186616.

Seall:1963:TVF

[228] Robert E. Seall. Truth-valued fluents and qualitative laws. Philosophy of Science, 30(1):36–40, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186617.

Ellson:1963:SCT

[229] D. G. Ellson. The scientists’ criterion of true observation. Philosophy of Science, 30(1):41–52, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186618.

King-Farlow:1963:TPN

[230] John King-Farlow. Truth preference and neuter propositions. Philosophy of Science, 30(1):53–59, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186619.

Henson:1963:MHS

[231] Richard G. Henson. Mr. Hanson on the symmetry of explanation and prediction. Philosophy of Science, 30(1):60–61, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186620.

Hamblin:1963:QAS

[232] C. L. Hamblin. Questions aren’t statements. Philosophy of Science, 30 (1):62–63, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186621.

Madden:1963:ENS

[233] Edward H. Madden. Ernest Nagel’s the structure of science. Philosophy of Science, 30(1):64–70, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186622.

Sweigart:1963:COC

[234] John Sweigart. Carnap’s ontology and challenges. Philosophy of Science, 30(1):71–80, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186623.

Beardsley:1963:BRB

[235] Monroe C. Beardsley. Book review: Logic: The Theory of Formal Inference by Alice Ambrose; Morris Lazerowitz. Philosophy of Science, 30(1): 81, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186624.

Lenzen:1963:BRB

[236] V. F. Lenzen. Book review: The Philosophical Impact of Contemporary Physics by Milic Capek. Philosophy of Science, 30(1):81–83, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186625.

Caton:1963:BRB

[237] Charles E. Caton. Book review: On the Nature of Meanings, a Philosophical Analysis by Niels Egmont Christensen. Philosophy of Science, 30(1):83–84, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186626.

Beer:1963:BRB

[238] Stafford Beer. Book review: Prediction and Optimal Decision: Philosophical Issues of a Science of Values by C. West Churchman. Philosophy of Science, 30(1):84–89, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186627.

Wolff:1963:BRB

[239] Kurt H. Wolff. Book review: Philosophy, Science and the Sociology of Knowledge by Irving Louis Horowitz. Philosophy of Science, 30(1):89–93, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186628.

Stannard:1963:BRBa

[240] Jerry Stannard. Book review: Science since Babylon by Derek J. De Solla Price. Philosophy of Science, 30(1):93–94, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186629.

Anonymous:1963:AIa

[241] Anonymous. Abstracts from inquiry. Philosophy of Science, 30(1):95, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186630.

Anonymous:1963:BMa

[242] Anonymous. Back matter. Philosophy of Science, 30(1):96–105, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186631.

Anonymous:1963:FMa

[243] Anonymous. Front matter. Philosophy of Science, 30(1):??, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186614.

Hanson:1963:LIP

[244] Norwood Russell Hanson. The law of inertia: A Philosopher’s touchstone. Philosophy of Science, 30(2):107–121, April 1963. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186419.

Martin:1963:PPP

[245] R. M. Martin. Performance, purpose, and permission. Philosophy of Science, 30(2):122–137, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186420.

Margenau:1963:MQSb

[246] Henry Margenau. Measurements and quantum states: Part II. Philosophy of Science, 30(2):138–157, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186421.

Caws:1963:SCC

[247] Peter Caws. Science, computers, and the complexity of nature. Philosophy of Science, 30(2):158–164, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186422.

Baumer:1963:WP

[248] W. H. Baumer. Von Wright’s paradoxes. Philosophy of Science, 30(2): 165–172, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186423.

Neidorf:1963:EP

[249] Robert Neidorf. Is Einstein a positivist? Philosophy of Science, 30(2):173– 188, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186424.

Vincent:1963:CAC

[250] R. H. Vincent. Concerning an alleged contradiction. Philosophy of Science, 30(2):189–194, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186425.

Parsons:1963:BRBa

[251] Howard L. Parsons. Book review: Perspectives in Philosophy: A Book of Readings by Robert N. Beck. Philosophy of Science, 30(2):195–196, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186426.

Dickie:1963:BRBa

[252] George Dickie. Book review: The Ascent of Life: A Philosophical Study of the Theory of Evolution by T. A. Goudge. Philosophy of Science, 30(2): 196–197, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186427.

Dickie:1963:BRBb

[253] George Dickie. Book review: Contemporary Theories of Knowledge by Thomas English Hill. Philosophy of Science, 30(2):197–198, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186428.

Goldstein:1963:BRBa

[254] Leon J. Goldstein. Book review: Toward a Science of Man in Society: A Positive Approach to the Integration of Social Knowledge by K. William Kapp. Philosophy of Science, 30(2):198–200, April 1963. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186429.

Anonymous:1963:AIb

[255] Anonymous. Abstracts from inquiry. Philosophy of Science, 30(2):201– 203, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186430.

Anonymous:1963:BMb

[256] Anonymous. Back matter. Philosophy of Science, 30(2):204–205, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186431.

Anonymous:1963:FMb

[257] Anonymous. Front matter. Philosophy of Science, 30(2):??, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186418.

Achinstein:1963:VAC

[258] Peter Achinstein. Variety and analogy in confirmation theory. Philosophy of Science, 30(3):207–221, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186233.

Carnap:1963:VAP

[259] Rudolf Carnap. Variety, analogy, and periodicity in inductive logic. Philosophy of Science, 30(3):222–227, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186234.

Ackermann:1963:NPC

[260] Robert Ackermann. A neglected proposal concerning simplicity. Philosophy of Science, 30(3):228–235, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186235.

Gellner:1963:NSS

[261] Ernest Gellner. Nature and society in social anthropology. Philosophy of Science, 30(3):236–251, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186236.

Salmon:1963:VI

[262] Wesley C. Salmon. On vindicating induction. Philosophy of Science, 30 (3):252–261, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186237.

Kennedy:1963:MPG

[263] Hubert C. Kennedy. The mathematical philosophy of giuseppe Peano. Philosophy of Science, 30(3):262–266, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186238.

Harrah:1963:MAI

[264] David Harrah. A model for applying information and utility functions. Philosophy of Science, 30(3):267–273, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186239.

Johnson:1963:CRB

[265] Rochelle J. Johnson. A commentary on “radical behaviorism”. Philosophy of Science, 30(3):274–285, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186240.

Kim:1963:LCD

[266] Jaegwon Kim. On the logical conditions of deductive explanation. Philosophy of Science, 30(3):286–291, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186241.

Goldstein:1963:BRBb

[267] Leon J. Goldstein. Book reviews: Man’s Place in Nature by Max Scheler; Hans Meyerhoff; Ressentiment by Max Scheler; Lewis Coser; William W. Holdheim. Philosophy of Science, 30(3):292–293, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186242.

Carney:1963:BRB

[268] James D. Carney. Book review: Generalization in Ethics by Marcus George Singer. Philosophy of Science, 30(3):293–295, July 1963. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186243.

Parsons:1963:BRBb

[269] Howard L. Parsons. Book review: The Nature of Philosophy: An Introduction by Samuel M. Thompson. Philosophy of Science, 30(3):295–297, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186244.

Sesonske:1963:BRB

[270] Alexander Sesonske. Book review: The Language of Ethics by Carl Wellman. Philosophy of Science, 30(3):298–299, July 1963. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186245.

Black:1963:BRB

[271] Max Black. Book review: Collected Papers of Charles Sanders Peirce by Arthur W. Burks. Philosophy of Science, 30(3):299–300, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186246.

Gluck:1963:BRBa

[272] Samuel E. Gluck. Book review: Growth in Living Systems by M. X. Zarrow. Philosophy of Science, 30(3):300–301, July 1963. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186247.

Anonymous:1963:BMc

[273] Anonymous. Back matter. Philosophy of Science, 30(3):302–303, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186248.

Anonymous:1963:FMc

[274] Anonymous. Front matter. Philosophy of Science, 30(3):??, July 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186232.

Myhill:1963:RLP

[275] John Myhill. Remarks on the language of physics. Philosophy of Science, 30(4):305–306, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186062.

Bohnert:1963:RMR

[276] H. G. Bohnert. Remarks on Myhill’s remarks on coordinate languages. Philosophy of Science, 30(4):307–308, October 1963. CODEN PHSCA6.

ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/186063.

Brodbeck:1963:MA

[277] May Brodbeck. Meaning and action. Philosophy of Science, 30(4):309– 324, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186064.

Rescher:1963:DSS

[278] Nicholas Rescher. Discrete state systems, Markov chains, and problems in the theory of scientific explanation and prediction. Philosophy of Science, 30(4):325–345, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186065.

Bunge:1963:GBB

[279] Mario Bunge. A general black box theory. Philosophy of Science, 30 (4):346–358, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186066.

Buck:1963:RP

[280] Roger C. Buck. Reflexive predictions. Philosophy of Science, 30(4):359– 369, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186067.

Grunbaum:1963:CPR

[281] Adolf Grünbaum. Comments on Professor Roger Buck’s paper “Reflexive Predictions”. Philosophy of Science, 30(4):370–372, October 1963. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186068.

Buck:1963:RG

[282] Roger C. Buck. Rejoinder to Grünbaum. Philosophy of Science, 30(4): 373–374, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186069.

Sacksteder:1963:DBS

[283] William Sacksteder. Diversity in the behavioral sciences. Philosophy of Science, 30(4):375–395, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186070.

Barker:1963:BRBa

[284] S. F. Barker. Book review: Classics in Biology: A Course of Selected Reading. Philosophy of Science, 30(4):396, October 1963. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186071.

Hinshaw:1963:BRB

[285] Virgil Hinshaw, Jr. Book review: The Origins of Scientific Thought by Giorgio de Santillana. Philosophy of Science, 30(4):396–398, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186072.

Haworth:1963:BRBa

[286] Lawrence Haworth. Book review: Reason in Society by Paul Diesing. Philosophy of Science, 30(4):398–399, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186073.

Goldstein:1963:BRBc

[287] Leon J. Goldstein. Book review: Values and Intentions, a Study in Value-Theory and Philosophy of Mind by J. N. Findlay. Philosophy of Science, 30(4):399–401, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186074.

Stannard:1963:BRBb

[288] Jerry Stannard. Book review: The Two Netherlanders: Humphrey Bradley and Cornelis Drebbel by L. E. Harris. Philosophy of Science, 30(4):401– 402, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186075.

MacKinnon:1963:BRB

[289] William J. MacKinnon. Book review: The Humanist Frame by Julian Huxley. Philosophy of Science, 30(4):402–403, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186076.

Haworth:1963:BRBb

[290] Lawrence Haworth. Book review: Society, Law, and Morality: Readings in Social Philosophy by Frederick A. Olafson. Philosophy of Science, 30 (4):403–404, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186077.

Goldstein:1963:BRBd

[291] Leon J. Goldstein. Book review: Dimensions of Freedom, an Analysis by Felix E. Oppenheim. Philosophy of Science, 30(4):404–405, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186078.

Linden:1963:BRB

[292] George W. Linden. Book review: Man’s New Image of Man by Oliver Reiser. Philosophy of Science, 30(4):405–406, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186079.

Barker:1963:BRBb

[293] S. F. Barker. Book review: Error and Deception in Science: Essays on Biological Aspects of Life by Jean Rostand; A. J. Pomerans. Philosophy of Science, 30(4):406–407, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186080.

Goldstein:1963:BRBe

[294] Leon J. Goldstein. Book review: Relativism and the Study of Man by Helmut Schoeck; James W. Wiggins. Philosophy of Science, 30(4):407– 408, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186081.

Gluck:1963:BRBb

[295] Samuel E. Gluck. Book review: The Measurement of Values by L. L. Thurstone. Philosophy of Science, 30(4):408–409, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186082.

Anonymous:1963:VI

[296] Anonymous. Volume information. Philosophy of Science, 30(4):411–415, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186060.

Anonymous:1963:BMd

[297] Anonymous. Back matter. Philosophy of Science, 30(4):??, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186083.

Anonymous:1963:FMd

[298] Anonymous. Front matter. Philosophy of Science, 30(4):??, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186061.

Putnam:1964:CCC

[299] Hilary Putnam. Comments on comments on comments: A reply to Margenau and Wigner. Philosophy of Science, 31(1):1–6, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186740.

Margenau:1964:RPP

[300] Henry Margenau and Eugene P. Wigner. Reply to Professor Putnam. Philosophy of Science, 31(1):7–9, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186741.

Losee:1964:UPA

[301] John Losee. The use of philosophical arguments in quantum physics. Philosophy of Science, 31(1):10–17, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186742.

Edel:1964:CUS

[302] Abraham Edel. The concept of the unconscious: Some analytic preliminaries. Philosophy of Science, 31(1):18–33, January 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186743.

Wilson:1964:AA

[303] P. R. Wilson. On the argument by analogy. Philosophy of Science, 31 (1):34–39, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186744.

Rothstein:1964:TSU

[304] Jerome Rothstein. Thermodynamics and some undecidable physical questions. Philosophy of Science, 31(1):40–48, January 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186745.

Pap:1964:TD

[305] Arthur Pap. Theory of definition. Philosophy of Science, 31(1):49–54, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186746.

Diamond:1964:RPT

[306] R. J. Diamond. Resolution of the paradox of tristram shandy. Philosophy of Science, 31(1):55–58, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186747.

Silvers:1964:SCQ

[307] Stuart Silvers. Some comments on Quine’s analysis of simplicity. Philosophy of Science, 31(1):59–61, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186748.

Skyrms:1964:PGT

[308] Brian Skyrms. Professor Grünbaum and teleological mechanisms. Philosophy of Science, 31(1):62–64, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186749.

Schlesinger:1964:FES

[309] G. Schlesinger. The formalization of empirical significance. Philosophy of Science, 31(1):65–67, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186750.

Madden:1964:MSM

[310] Edward H. Madden and Murray J. Kiteley. Mr. Schlesinger on the M-K theory. Philosophy of Science, 31(1):68–70, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186751.

Riepe:1964:ZSO

[311] Dale Riepe. Zen and the scientific outlook. Philosophy of Science, 31 (1):71–74, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186752.

Rieser:1964:ICM

[312] Max Rieser. 1961 International Colloquy for Methodology of Sciences in Warsaw, Poland, Sept. 18-23, 1961. Philosophy of Science, 31(1):75–81, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186753.

Anonymous:1964:AIa

[313] Anonymous. Abstracts from inquiry. Philosophy of Science, 31(1):82–86, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186754.

Anonymous:1964:ABJ

[314] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 31(1):87–90, January 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186755.

Anonymous:1964:BMa

[315] Anonymous. Back matter. Philosophy of Science, 31(1):91–100, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186756.

Anonymous:1964:FMa

[316] Anonymous. Front matter. Philosophy of Science, 31(1):??, January 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186739.

Todd:1964:CCP

[317] William Todd. Counterfactual conditionals and the presuppositions of induction. Philosophy of Science, 31(2):101–110, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185987.

Shapere:1964:CES

[318] Dudley Shapere. The causal efficacy of space. Philosophy of Science, 31(2):111–121, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185988.

Martin:1964:EVU

[319] Michael Martin. The explanatory value of the unconscious. Philosophy of Science, 31(2):122–132, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185989.

Dretske:1964:PR

[320] Fred I. Dretske. Particular reidentification. Philosophy of Science, 31(2): 133–142, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185990.

Suppes:1964:EUH

[321] Patrick Suppes. On an example of unpredictability in human behavior. Philosophy of Science, 31(2):143–148, April 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185991.

Walsh:1964:SWC

[322] Vivian Charles Walsh. The status of Welfare comparisons. Philosophy of Science, 31(2):149–155, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185992.

Fine:1964:PGP

[323] Arthur Fine. Physical geometry and physical laws. Philosophy of Science, 31(2):156–162, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185993.

Simons:1964:PNA

[324] Leo Simons. Professor Nagel on abstractive theories and experimental laws. Philosophy of Science, 31(2):163–167, April 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185994.

Copeland:1964:PDS

[325] John W. Copeland. Philosophy disguised as science. Philosophy of Science, 31(2):168–172, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185995.

Hurrell:1964:ITT

[326] Paul Manson Hurrell. Interrogatives, testability and truth-value. Philosophy of Science, 31(2):173–182, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185996.

Hinshaw:1964:BRBa

[327] Virgil Hinshaw, Jr. Book review: Intuition and Science by Mario Bunge. Philosophy of Science, 31(2):183–184, April 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185997.

Scriven:1964:BRB

[328] Michael Scriven. Book review: Science in Authority by Lancelot Hogben. Philosophy of Science, 31(2):184–186, April 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185998.

Martin:1964:BRB

[329] Michael Martin. Book review: Freud’s Concept of Repression and Defense, Its Theoretical and Observational Language by Peter Madison. Philosophy of Science, 31(2):186–188, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185999.

Burgener:1964:BRB

[330] R. J. C. Burgener. Book review: Free Will and Determinism by Allan M. Munn. Philosophy of Science, 31(2):188–189, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186000.

Goldstein:1964:BRB

[331] Leon J. Goldstein. Book review: An Inquiry into the Freedom of Decision by Harald Ofstad. Philosophy of Science, 31(2):189–190, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186001.

Hinshaw:1964:BRBb

[332] Virgil Hinshaw, Jr. Book review: Thought and Language by L. S. Vygotsky; E. Hanfmann; G. Vakar. Philosophy of Science, 31(2):190–191, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186002.

Suits:1964:BRB

[333] Bernard Suits. Book review: The Social Theories of Talcott Parsons by Max Black; Alfred L. Baldwin; Urie Bronfenbrenner; Edward C. Devereux; Andrew Hacker; Henry A. Landsberger; Chandler Morse; Talcott Parsons; William Foote Whyte; Robin M. Williams. Philosophy of Science, 31(2): 192–195, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186003.

Anonymous:1964:BMb

[334] Anonymous. Back matter. Philosophy of Science, 31(2):196, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186004.

Anonymous:1964:FMb

[335] Anonymous. Front matter. Philosophy of Science, 31(2):??, April 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185986.

Sellars:1964:IV

[336] Wilfrid Sellars. Induction as vindication. Philosophy of Science, 31(3): 197–231, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186085.

Rosen:1964:GPD

[337] Robert Rosen. The Gibbs’ Paradox and the distinguishability of physical systems. Philosophy of Science, 31(3):232–236, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186086.

Krause:1964:DRP

[338] Merton S. Krause. Disconfirmative results and prior commitments. Philosophy of Science, 31(3):237–240, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186087.

Workman:1964:WME

[339] Rollin W. Workman. What makes an explanation. Philosophy of Science, 31(3):241–254, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186088.

Rogers:1964:MPA

[340] Robert Rogers. Mathematical and philosophical analyses. Philosophy of Science, 31(3):255–264, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186089.

Berenda:1964:CIP

[341] Carlton W. Berenda. On the cosmological indeterminacy principle of Mc-Crae. Philosophy of Science, 31(3):265–270, July 1964. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186090.

Fleming:1964:SQE

[342] John J. Fleming. Sub-quantum entities. Philosophy of Science, 31(3):271– 274, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186091.

Horowitz:1964:PDT

[343] Irving Louis Horowitz. Professionalism and disciplinarianism: Two styles of sociological performance. Philosophy of Science, 31(3):275–281, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186092.

Rozeboom:1964:SOS

[344] William W. Rozeboom. Of selection operators and semanticists. Philosophy of Science, 31(3):282–285, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186093.

Barr:1964:ECP

[345] H. J. Barr. The epistemology of causality from the point of view of evolutionary biology. Philosophy of Science, 31(3):286–288, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186094.

Frings:1964:RMM

[346] Manfred S. Frings. A reply to Mr. Mattick’s article on Marxism and the New Physics. Philosophy of Science, 31(3):289–293, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186095. See [206].

Barnes:1964:PSF

[347] J. A. Barnes. Physical and social facts in anthropology. Philosophy of Science, 31(3):294–297, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186096.

Anonymous:1964:BMc

[348] Anonymous. Back matter. Philosophy of Science, 31(3):298–299, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186097.

Anonymous:1964:FMc

[349] Anonymous. Front matter. Philosophy of Science, 31(3):??, July 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186084.

Hanson:1964:SPC

[350] Norwood Russell Hanson. Stability proofs and consistency proofs: A loose analogy. Philosophy of Science, 31(4):301–318, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186261.

Hesse:1964:ACT

[351] Mary Hesse. Analogy and confirmation theory. Philosophy of Science, 31 (4):319–327, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186262.

Achinstein:1964:MAT

[352] Peter Achinstein. Models, analogies, and theories. Philosophy of Science, 31(4):328–350, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186263.

Agassi:1964:AG

[353] Joseph Agassi. Analogies as generalizations. Philosophy of Science, 31 (4):351–356, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186264.

Ellis:1964:ND

[354] Brian Ellis. On the nature of dimensions. Philosophy of Science, 31(4): 357–380, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186265.

Anonymous:1964:AIb

[355] Anonymous. Abstracts from inquiry. Philosophy of Science, 31(4):381– 383, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186266.

Anonymous:1964:VI

[356] Anonymous. Volume information. Philosophy of Science, 31(4):385–389, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186259.

Anonymous:1964:BMd

[357] Anonymous. Back matter. Philosophy of Science, 31(4):390, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186267.

Anonymous:1964:FMd

[358] Anonymous. Front matter. Philosophy of Science, 31(4):??, October 1964. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186260.

Lehman:1965:FEB

[359] Hugh Lehman. Functional explanation in biology. Philosophy of Science, 32(1):1–20, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186706.

Anonymous:1965:RRE

[360] Anonymous. Results of recent elections for officers and members of the governing board for the philosophy of science association. Philosophy of Science, 32(1):20, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186707.

Schlegel:1965:PIM

[361] Richard Schlegel. The problem of infinite matter in steady-state cosmology. Philosophy of Science, 32(1):21–31, January 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186708.

Vickers:1965:SRC

[362] John M. Vickers. Some remarks on coherence and subjective probability. Philosophy of Science, 32(1):32–38, January 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186709.

Graves:1965:MMR

[363] John C. Graves and James E. Roper. Measuring measuring rods. Philosophy of Science, 32(1):39–56, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186710.

Raab:1965:MM

[364] Francis V. Raab. Of minds and molecules. Philosophy of Science, 32(1): 57–72, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186711.

Brender:1965:HPC

[365] Myron Brender. Hospers on psychoanalysis: A critique. Philosophy of Science, 32(1):73–83, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186712.

Anonymous:1965:AIa

[366] Anonymous. Abstracts from inquiry. Philosophy of Science, 32(1):84–86, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186713.

Anonymous:1965:AD

[367] Anonymous. Abstracts from dialectica. Philosophy of Science, 32(1):87– 91, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186714.

Anonymous:1965:BMa

[368] Anonymous. Back matter. Philosophy of Science, 32(1):92–103, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186715.

Anonymous:1965:FMa

[369] Anonymous. Front matter. Philosophy of Science, 32(1):??, January 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186705.

Mellor:1965:EED

[370] D. H. Mellor. Experimental error and deducibility. Philosophy of Science, 32(2):105–122, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186404.

Hochberg:1965:BBP

[371] Herbert Hochberg. On being and being presented. Philosophy of Science, 32(2):123–136, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186405.

Salmon:1965:SPP

[372] Wesley C. Salmon. The status of prior probabilities in statistical explanation. Philosophy of Science, 32(2):137–146, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186406.

Kyburg:1965:SP

[373] Henry E. Kyburg, Jr. Salmon’s paper. Philosophy of Science, 32(2):147– 151, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186407.

Salmon:1965:RK

[374] Wesley C. Salmon. Reply to Kyburg. Philosophy of Science, 32(2):152– 154, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186408.

Ackermann:1965:DSE

[375] Robert Ackermann. Deductive scientific explanation. Philosophy of Science, 32(2):155–167, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186409.

Newman:1965:ES

[376] Fred Newman. Explanation sketches. Philosophy of Science, 32(2):168– 172, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186410.

Grunberg:1965:RP

[377] Emile Grunberg and Franco Modigliani. Reflexive prediction. Philosophy of Science, 32(2):173–174, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186411.

Butts:1965:WRW

[378] Robert E. Butts. On Walsh’s reading of Whewell’s view of necessity. Philosophy of Science, 32(2):175–181, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186412.

Lamb:1965:TSS

[379] Matthew L. Lamb. Towards a synthetization of the sciences. Philosophy of Science, 32(2):182–191, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186413.

Lande:1965:SGE

[380] Alfred Landé. Solution of the Gibbs entropy paradox. Philosophy of Science, 32(2):192–193, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186414.

Madden:1965:GBP

[381] Edward H. Madden. E. G. Boring’s philosophy of science. Philosophy of Science, 32(2):194–201, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186415.

Kurtz:1965:BRB

[382] Paul Kurtz. Book review: Philosophical Problems of Psychology by Edward H. Madden. Philosophy of Science, 32(2):202–203, April 1965. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186416.

Anonymous:1965:BMb

[383] Anonymous. Back matter. Philosophy of Science, 32(2):204, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186417.

Anonymous:1965:FMb

[384] Anonymous. Front matter. Philosophy of Science, 32(2):??, April 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186403.

Adams:1965:ETI

[385] Ernest W. Adams. Elements of a theory of inexact measurement. Philosophy of Science, 32(3/4):205–228, July/October 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186515.

Robinson:1965:MST

[386] Richard E. Robinson. Measurement and statistics: Towards a clarification of the theory of “permissible statistics”. Philosophy of Science, 32(3/4): 229–243, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186516.

Manier:1965:TEP

[387] Edward Manier. The theory of evolution as personal knowledge. Philosophy of Science, 32(3/4):244–252, July/October 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186517. See reply [430].

Skyrms:1965:FVI

[388] Brian Skyrms. On failing to vindicate induction. Philosophy of Science, 32(3/4):253–268, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186518.

Hacking:1965:SV

[389] Ian Hacking. Salmon’s vindication. Philosophy of Science, 32(3/4):269– 271, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186519.

Clatterbaugh:1965:GOP

[390] Kenneth C. Clatterbaugh. General ontology and the principle of acquaintance. Philosophy of Science, 32(3/4):272–276, July/October 1965. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186520.

Allaire:1965:OAR

[391] Edwin B. Allaire. Ontology and acquaintance: A reply to Clatterbaugh. Philosophy of Science, 32(3/4):277–280, July/October 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/186521.

Sinisi:1965:DCN

[392] Vito F. Sinisi. Discussion: “∈” and common names. Philosophy of Science, 32(3/4):281–286, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186522.

Seaman:1965:DD

[393] Francis Seaman. In defense of Duhem. Philosophy of Science, 32(3/4): 287–294, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186523.

Laudan:1965:GBD

[394] Laurens Laudan. Grünbaum on “The Duhemian Argument”. Philosophy of Science, 32(3/4):295–299, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186524. See [6].

Luce:1965:FAM

[395] R. Duncan Luce. A “fundamental” axiomatization of multiplicative power relations among three variables. Philosophy of Science, 32(3/4):301–309, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186525.

Tredwell:1965:PC

[396] R. F. Tredwell. The problem of counterfactuals. Philosophy of Science, 32(3/4):310–323, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186526.

Winnie:1965:TTP

[397] John A. Winnie. Theoretical terms and partial definitions. Philosophy of Science, 32(3/4):324–328, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186527.

Resnik:1965:FTI

[398] Michael David Resnik. Frege’s theory of incomplete entities. Philosophy of Science, 32(3/4):329–341, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186528.

Massey:1965:FSS

[399] Gerald J. Massey. Four simple systems of modal propositional logic. Philosophy of Science, 32(3/4):342–355, July/October 1965. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186529.

Politis:1965:LF

[400] Constantine Politis. Limitations of formalization. Philosophy of Science, 32(3/4):356–360, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186530.

Barrett:1965:QSL

[401] Robert Barrett. Quine, synonymy and logical truth. Philosophy of Science, 32(3/4):361–367, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186531.

Jones:1965:NRK

[402] Robert M. Jones. The non-reducibility of Koopman’s theorems of probability in Carnap’s system for MC. Philosophy of Science, 32(3/4):368–369, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186532.

Clendinnen:1965:KVI

[403] F. John Clendinnen. Katz on the vindication of induction. Philosophy of Science, 32(3/4):370–376, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186533.

Kahane:1965:NGE

[404] Howard Kahane. Nelson Goodman’s entrenchment theory. Philosophy of Science, 32(3/4):377–383, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186534.

Anonymous:1965:ABJ

[405] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 32(3/4):384–387, July/October 1965. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186535.

Anonymous:1965:AIb

[406] Anonymous. Abstracts from inquiry. Philosophy of Science, 32(3/4):388– 390, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186536.

Anonymous:1965:VI

[407] Anonymous. Volume information. Philosophy of Science, 32(3/4):391– 394, July/October 1965. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186513.

Anonymous:1965:BMc

[408] Anonymous. Back matter. Philosophy of Science, 32(3/4):395–396, July/ October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186537.

Anonymous:1965:FMc

[409] Anonymous. Front matter. Philosophy of Science, 32(3/4):??, July/ October 1965. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186514.

Martin:1966:TCR

[410] R. M. Martin. On theoretical constructs and Ramsey constants. Philosophy of Science, 33(1/2):1–13, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186433.

Suppes:1966:PAN

[411] Patrick Suppes. The probabilistic argument for a non-classical logic of quantum mechanics. Philosophy of Science, 33(1/2):14–21, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186434.

Sneed:1966:NAP

[412] Joseph D. Sneed. Von Neumann’s argument for the projection postulate. Philosophy of Science, 33(1/2):22–39, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186435.

Mischel:1966:PAE

[413] Theodore Mischel. Pragmatic aspects of explanation. Philosophy of Science, 33(1/2):40–60, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186436.

Leach:1966:DRE

[414] James Leach. Dray on rational explanation. Philosophy of Science, 33(1/2):61–69, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186437.

Ackermann:1966:PU

[415] Robert J. Ackermann. Projecting unprojectibles. Philosophy of Science, 33(1/2):70–75, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186438.

Edwards:1966:TFD

[416] Rem B. Edwards. The truth and falsity of definitions. Philosophy of Science, 33(1/2):76–79, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186439.

Stenner:1966:BRB

[417] Alfred Stenner. Book review: Anatomy of Inquiry by Israel Scheffler. Philosophy of Science, 33(1/2):80–82, March/June 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186440.

Madden:1966:BRB

[418] Edward H. Madden. Book review: The Anatomy of Inquiry by Israel Scheffler. Philosophy of Science, 33(1/2):82–84, March/June 1966. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186441.

Rose:1966:BRBa

[419] Alan Rose. Book review: Formal Methods. An Introduction to Symbolic Logic and to the Study of Effective Operations in Arithmetic and Logic by Evert W. Beth. Philosophy of Science, 33(1/2):84–85, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186442.

Rose:1966:BRBb

[420] Allan Rose. Book review: Logico-Philosophical Studies by Albert Menne; Horace S. Glover. Philosophy of Science, 33(1/2):85, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186443.

Fisk:1966:BRB

[421] Milton Fisk. Book review: Reasoning and Logic by Richard B. Angell; Sterling P. Lamprecht. Philosophy of Science, 33(1/2):85–87, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186444.

Anonymous:1966:ADa

[422] Anonymous. Abstracts from dialectica. Philosophy of Science, 33(1/2): 88–89, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186445.

Anonymous:1966:BMa

[423] Anonymous. Back matter. Philosophy of Science, 33(1/2):90–198, March/ June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186460.

Addis:1966:FMP

[424] Laird Addis. Freedom and the Marxist philosophy of history. Philosophy of Science, 33(1/2):101–117, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186446.

Margolis:1966:OI

[425] Joseph Margolis. Objectivism and interactionism. Philosophy of Science, 33(1/2):118–123, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186447.

Diesing:1966:OVS

[426] Paul Diesing. Objectivism vs. subjectivism in the social sciences. Philosophy of Science, 33(1/2):124–133, March/June 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186448.

Cannavo:1966:ERP

[427] S. Cannavo. Extensionality and randomness in probability sequences. Philosophy of Science, 33(1/2):134–146, March/June 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186449.

Ghiselin:1966:SPB

[428] Michael T. Ghiselin. On semantic pitfalls of biological adaptation. Philosophy of Science, 33(1/2):147–153, March/June 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186450.

Axinn:1966:FSR

[429] Sidney Axinn. Fallacy of the single risk. Philosophy of Science, 33(1/2): 154–162, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186451.

Grene:1966:MMB

[430] Marjorie Grene. Mr. Manier’s “Theory of Evolution as Personal Knowledge”: A quasi-reply. Philosophy of Science, 33(1/2):163–164, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186452. See [387].

Beard:1966:DPC

[431] Robert W. Beard. Deduction, prediction and completeness conditions. Philosophy of Science, 33(1/2):165–167, March/June 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186453.

Ackermann:1966:CME

[432] Robert Ackermann and Alfred Stenner. A corrected model of explanation. Philosophy of Science, 33(1/2):168–171, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186454.

Watson:1966:GDC

[433] Richard A. Watson. Is geology different: A critical discussion of “The Fabric of Geology”. Philosophy of Science, 33(1/2):172–185, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186455.

Bricklin:1966:BRB

[434] Barry Bricklin. Book review: The Philosophy of Personality by Radhakamal Mukerjee. Philosophy of Science, 33(1/2):186–189, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186456.

Murphree:1966:BRB

[435] Idas Murphree. Book review: The Pragmatism of C. S. Peirce. An Analytical Study by Hjalmar Wennerberg. Philosophy of Science, 33(1/2): 189–190, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186457.

Parsons:1966:BRB

[436] Howard L. Parsons. Book review: Search for a Method by Jean-Paul Sartre; Hazel E. Barnes. Philosophy of Science, 33(1/2):190–192, March/ June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186458.

Anonymous:1966:AIa

[437] Anonymous. Abstracts from inquiry. Philosophy of Science, 33(1/2): 193–196, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186459.

Anonymous:1966:FMa

[438] Anonymous. Front matter. Philosophy of Science, 33(1/2):??, March/June 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186432.

Kouznetsov:1966:CR

[439] B. Kouznetsov. Complementarity and relativity. Philosophy of Science, 33 (3):199–209, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186269.

Swinburne:1966:CH

[440] R. G. Swinburne. Cosmological horizons. Philosophy of Science, 33(3): 210–214, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186270.

Clendinnen:1966:IO

[441] F. John Clendinnen. Induction and objectivity. Philosophy of Science, 33 (3):215–229, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186271.

Skyrms:1966:NNP

[442] Brian Skyrms. Nomological necessity and the paradoxes of confirmation. Philosophy of Science, 33(3):230–249, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186272.

Kyburg:1966:PD

[443] H. E. Kyburg, Jr. Probability and decision. Philosophy of Science, 33(3): 250–261, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186273.

Emmerich:1966:SDF

[444] David S. Emmerich and James G. Greeno. Some decision factors in scientific investigation. Philosophy of Science, 33(3):262–270, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186274.

Klemke:1966:LL

[445] E. D. Klemke. The laws of logic. Philosophy of Science, 33(3):271–277, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186275.

Madden:1966:EPH

[446] Edward H. Madden. Explanation in psychoanalysis and history. Philosophy of Science, 33(3):278–286, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186276.

Brodbeck:1966:OIR

[447] May Brodbeck. Objectivism and interaction: A reaction to Margolis. Philosophy of Science, 33(3):287–292, September 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186277.

Margolis:1966:RRS

[448] Joseph Margolis. Reply to a reaction: Second remarks on Brodbeck’s objectivism. Philosophy of Science, 33(3):293–300, September 1966. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186278.

Anonymous:1966:BMb

[449] Anonymous. Back matter. Philosophy of Science, 33(3):301–302, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186279.

Anonymous:1966:FMb

[450] Anonymous. Front matter. Philosophy of Science, 33(3):??, September 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186268.

Copeland:1966:MPE

[451] Arthur H. Copeland, Sr. Mathematical proof and experimental proof. Philosophy of Science, 33(4):303–316, December 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186634.

Cohen:1966:CQM

[452] Leon Cohen. Can quantum mechanics be formulated as a classical probability theory? Philosophy of Science, 33(4):317–322, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186635.

Simon:1966:CC

[453] Herbert A. Simon and Nicholas Rescher. Cause and counterfactual. Philosophy of Science, 33(4):323–340, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186636.

Rietdijk:1966:RPD

[454] C. W. Rietdijk. A rigorous proof of determinism derived from the Special Theory of Relativity. Philosophy of Science, 33(4):341–344, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186637.

Mellor:1966:IE

[455] D. H. Mellor. Inexactness and explanation. Philosophy of Science, 33 (4):345–359, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186638.

Kekes:1966:PIT

[456] John Kekes. Physicalism, the identity theory, and the doctrine of emergence. Philosophy of Science, 33(4):360–375, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186639.

Thompson:1966:CTD

[457] John W. Thompson. Coombs’ theory of data. Philosophy of Science, 33 (4):376–382, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186640.

Butts:1966:FPT

[458] Robert E. Butts. Feyerabend and the pragmatic theory of observation. Philosophy of Science, 33(4):383–394, December 1966. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186641.

Agassi:1966:MR

[459] Joseph Agassi. The mystery of the ravens. Philosophy of Science, 33(4): 395–402, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186642.

Margenau:1966:BRB

[460] Henry Margenau. Book review: Einstein und Die Sowjetphilosophie by S. Müller-Markus. Philosophy of Science, 33(4):403–404, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186643.

Anonymous:1966:ADb

[461] Anonymous. Abstracts from dialectica. Philosophy of Science, 33(4): 405–407, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186644.

Anonymous:1966:ABJ

[462] Anonymous. Abstracts from British journal for the philosophy of science. Philosophy of Science, 33(4):408–411, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186645.

Anonymous:1966:AIb

[463] Anonymous. Abstracts from inquiry. Philosophy of Science, 33(4):412– 414, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186646.

Anonymous:1966:EMS

[464] Anonymous. Erratum: Measurement and statistics: Towards a clarification of the theory of “permissible statistics”. Philosophy of Science, 33(4):414, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186647.

Anonymous:1966:VI

[465] Anonymous. Volume information. Philosophy of Science, 33(4):415–419, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186632.

Anonymous:1966:BMc

[466] Anonymous. Back matter. Philosophy of Science, 33(4):420–422, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186648.

Anonymous:1966:FMc

[467] Anonymous. Front matter. Philosophy of Science, 33(4):??, December 1966. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186633.

Mellor:1967:IE

[468] D. H. Mellor. Imprecision and explanation. Philosophy of Science, 34(1): 1–9, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186569.

Hochberg:1967:DP

[469] Herbert Hochberg. Dispositional properties. Philosophy of Science, 34(1): 10–17, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186570.

Michalos:1967:PRP

[470] Alex C. Michalos. Postulates of rational preference. Philosophy of Science, 34(1):18–22, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186571.

Goldstein:1967:TH

[471] Leon J. Goldstein. Theory in history. Philosophy of Science, 34(1):23– 40, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186572.

Suchting:1967:DEP

[472] W. A. Suchting. Deductive explanation and prediction revisited. Philosophy of Science, 34(1):41–52, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186573.

Purtill:1967:KSR

[473] Richard L. Purtill. Kuhn on scientific revolutions. Philosophy of Science, 34(1):53–58, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186574.

Swanson:1967:DT

[474] J. W. Swanson. On the D-thesis. Philosophy of Science, 34(1):59–68, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186575.

Narveson:1967:BRBa

[475] Anne Narveson. Book review: A Companion to Wittgenstein’s Tractatus by Max Black. Philosophy of Science, 34(1):69–73, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186576.

Narveson:1967:BRBb

[476] Jan Narveson. Book review: Ethics by William Frankena. Philosophy of Science, 34(1):74, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186577.

Narveson:1967:BRBc

[477] Jan Narveson. Book review: Ethics and Metaethics by Raziel Abelson. Philosophy of Science, 34(1):74–75, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186578.

Handy:1967:BRB

[478] Rollo Handy. Book review: Explanation in Social Science by Robert Brown. Philosophy of Science, 34(1):75–76, March 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186579.

Thomas:1967:BRBa

[479] Ivo Thomas. Book review: Fundamentals of Logic by James D. Carney; Richard K. Scheer. Philosophy of Science, 34(1):76–77, March 1967. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186580.

Caton:1967:BRB

[480] Charles E. Caton. Book review: Knowledge and Experience, Proceedings of the 1962 Oberlin Colloquium in Philosophy by C. D. Rollins. Philosophy of Science, 34(1):77–79, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186581.

Thomas:1967:BRBb

[481] Ivo Thomas. Book review: The Logical Systems of Lesniewski by Eugene C. Luschei. Philosophy of Science, 34(1):79–80, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186582.

Narveson:1967:BRBd

[482] Anne Narveson. Book review: The Philosophy of Wittgenstein by George Pitcher. Philosophy of Science, 34(1):80–83, March 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186583.

Harrison:1967:BRB

[483] Frank R. Harrison, III. Book review: A Portrait of Aristotle by Marjorie Grene. Philosophy of Science, 34(1):83–84, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186584.

Yolton:1967:BRB

[484] John W. Yolton. Book review: The Primary World of Senses; A Vindication of Sensory Experience by Erwin Straus; Jacob Needleman. Philosophy of Science, 34(1):84–85, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186585.

Anonymous:1967:ASa

[485] Anonymous. Abstracts from synthese. Philosophy of Science, 34(1):86– 87, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186586.

Anonymous:1967:BMa

[486] Anonymous. Back matter. Philosophy of Science, 34(1):88–104, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186587.

Anonymous:1967:FMa

[487] Anonymous. Front matter. Philosophy of Science, 34(1):??, March 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186568.

Meehl:1967:TTP

[488] Paul E. Meehl. Theory-testing in psychology and physics: A methodological paradox. Philosophy of Science, 34(2):103–115, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186099.

Ellis:1967:CDS

[489] Brian Ellis and Peter Bowman. Conventionality in distant simulataneity. Philosophy of Science, 34(2):116–136, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186100.

Schaffner:1967:AR

[490] Kenneth F. Schaffner. Approaches to reduction. Philosophy of Science, 34(2):137–147, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186101.

Suits:1967:WG

[491] Bernard Suits. What is a game? Philosophy of Science, 34(2):148–156, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186102.

Lehman:1967:BSR

[492] Hugh Lehman. Are biological species real? Philosophy of Science, 34(2): 157–167, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186103.

Bock:1967:TDM

[493] Philip K. Bock. Three descriptive models of social structure. Philosophy of Science, 34(2):168–174, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186104.

Butts:1967:PMW

[494] Robert E. Butts. Professor Marcucci on Whewell’s idealism. Philosophy of Science, 34(2):175–183, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186105.

Kahane:1967:RA

[495] Howard Kahane. Reply to Ackermann. Philosophy of Science, 34(2):184– 187, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186106.

Ackermann:1967:CD

[496] Robert J. Ackermann. Conflict and decision. Philosophy of Science, 34 (2):188–193, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186107.

Michalos:1967:BRBa

[497] Alex C. Michalos. Book review: The Character of Physical Law by Richard Feynman. Philosophy of Science, 34(2):194, June 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186108.

Michalos:1967:BRBb

[498] Alex C. Michalos. Book review: Decision and the Condition of Man by Paul Kurtz. Philosophy of Science, 34(2):194–195, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186109.

Michalos:1967:BRBc

[499] Alex C. Michalos. Book review: Statistical and Inductive Probabilities by Hugues Leblanc. Philosophy of Science, 34(2):195–196, June 1967.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186110.

Lajewski:1967:BRB

[500] C. Lajewski. Book review: Elements of Mathematical Logic by Jan Lukasiewicz; O. Wojtasiewicz. Philosophy of Science, 34(2):197–198, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186111.

Thomas:1967:BRBc

[501] Ivo Thomas. Book review: Galen and the Syllogism by Nicholas Rescher. Philosophy of Science, 34(2):198–200, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186112.

Resnik:1967:BRBa

[502] Michael David Resnik. Book review: Frege’s Logical Theory by Robert Sternfeld. Philosophy of Science, 34(2):201–202, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186113.

Silvers:1967:BRB

[503] Stuart Silvers. Book review: Choice and Chance: An Introduction to Inductive Logic by Brian Skyrms. Philosophy of Science, 34(2):202–203, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186114.

Anonymous:1967:AI

[504] Anonymous. Abstracts from inquiry. Philosophy of Science, 34(2):204– 208, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186115.

Anonymous:1967:ASb

[505] Anonymous. Abstracts from synthese. Philosophy of Science, 34(2):209– 210, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186116.

Anonymous:1967:BMb

[506] Anonymous. Back matter. Philosophy of Science, 34(2):??, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3857061.

Anonymous:1967:FMb

[507] Anonymous. Front matter. Philosophy of Science, 34(2):??, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186098.

Earman:1967:GBT

[508] John Earman. On going backward in time. Philosophy of Science, 34(3): 211–222, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186497.

Jarvie:1967:TFS

[509] I. C. Jarvie. On theories of fieldwork and the scientific character of social anthropology. Philosophy of Science, 34(3):223–242, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186498.

Walsh:1967:SCS

[510] Vivian Charles Walsh. On the significance of choice sets with incompatibilities. Philosophy of Science, 34(3):243–250, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186499.

Addis:1967:PA

[511] Laird Addis. Particulars and acquaintance. Philosophy of Science, 34(3): 251–259, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186500.

Allaire:1967:TRI

[512] Edwin B. Allaire. Things, relations and identity. Philosophy of Science, 34 (3):260–272, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186501.

Goh:1967:NES

[513] S. T. Goh. Newman and explanation-sketches. Philosophy of Science, 34 (3):273–275, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186502.

Angel:1967:EPP

[514] R. B. Angel. Explanation and prediction: A plea for reason. Philosophy of Science, 34(3):276–282, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186503.

Salmon:1967:BRB

[515] Wesley C. Salmon. Book review: Probability and the Logic of Rational Belief by Henry E. Kyburg. Philosophy of Science, 34(3):283–285, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186504.

Armour:1967:BRB

[516] Leslie Armour. Book review: Experience and Conceptual Activity by J. M. Burgers. Philosophy of Science, 34(3):286–287, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186505.

Grunstra:1967:BRB

[517] Bernard R. Grunstra. Book review: Basic Concepts of Measurement by Brian Ellis. Philosophy of Science, 34(3):288–291, September 1967. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186506.

Narveson:1967:BRBe

[518] Jan Narveson. Book review: Distributive Justice by Nicholas Rescher. Philosophy of Science, 34(3):291–294, September 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186507.

Thomas:1967:BRBd

[519] Ivo Thomas. Book review: Introduction to Logic by William of Sherwood; Norman Kretzmann. Philosophy of Science, 34(3):295–296, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186508.

Resnik:1967:BRBb

[520] Michael David Resnik. Book review: A Study of Frege by Jeremy D. B. Walker. Philosophy of Science, 34(3):296–298, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186509.

Anonymous:1967:ABJa

[521] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 34(3):299–302, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186510.

Anonymous:1967:ASc

[522] Anonymous. Abstracts from synthese. Philosophy of Science, 34(3):303– 304, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186511.

Anonymous:1967:BMc

[523] Anonymous. Back matter. Philosophy of Science, 34(3):??, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186512.

Anonymous:1967:FMc

[524] Anonymous. Front matter. Philosophy of Science, 34(3):??, September 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186496.

Savage:1967:DTP

[525] Leonard J. Savage. Difficulties in the theory of personal probability. Philosophy of Science, 34(4):305–310, December 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186119. See comments [527].

Hacking:1967:SMR

[526] Ian Hacking. Slightly more realistic personal probability. Philosophy of Science, 34(4):311–325, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186120.

Shimony:1967:APP

[527] Abner Shimony. Amplifying personal probability theory: Comments on L. J. Savage’s “Difficulties in the Theory of Personal Probability”. Philosophy of Science, 34(4):326–332, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186121. See [525].

Bolker:1967:SAU

[528] Ethan D. Bolker. A simultaneous axiomatization of utility and subjective probability. Philosophy of Science, 34(4):333–340, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186122.

Bohnert:1967:CRS

[529] Herbert G. Bohnert. Communication by Ramsey-sentence clause. Philosophy of Science, 34(4):341–347, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186123.

Krantz:1967:EMS

[530] David H. Krantz. Extensive measurement in semiorders. Philosophy of Science, 34(4):348–362, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186124.

Jobe:1967:SRW

[531] Evan K. Jobe. Some recent work on the problem of law. Philosophy of Science, 34(4):363–381, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186125.

Hendry:1967:BRB

[532] Herbert E. Hendry. Book review: Techniques of Deductive Inference by Hughes Leblanc. Philosophy of Science, 34(4):382–383, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186126.

Schlegel:1967:BRB

[533] Richard Schlegel. Book review: Symmetries and Reflections: Scientific Essays of Eugene P. Wigner edited by Walter J. Moore; Michael Scriven. Philosophy of Science, 34(4):383–386, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186127.

Madden:1967:BRB

[534] Edward H. Madden. Book review: Completeness in Science by Richard Schlegel. Philosophy of Science, 34(4):386–388, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186128.

Braybrooke:1967:BRB

[535] David Braybrooke. Book review: Analytical Philosophy of History by Arthur C. Danto. Philosophy of Science, 34(4):388–392, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186129.

Gorovitz:1967:BRB

[536] Samuel Gorovitz. Book review: Introduction to Logic by Nicholas Rescher. Philosophy of Science, 34(4):392–396, December 1967. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186130.

Anonymous:1967:ABJb

[537] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 34(4):397–398, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186131.

Anonymous:1967:AD

[538] Anonymous. Abstracts from dialectica. Philosophy of Science, 34(4): 399–400, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186132.

Anonymous:1967:ASd

[539] Anonymous. Abstracts from synthese. Philosophy of Science, 34(4):401– 403, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186133.

Anonymous:1967:VI

[540] Anonymous. Volume information. Philosophy of Science, 34(4):405–409, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186117.

Anonymous:1967:BMd

[541] Anonymous. Back matter. Philosophy of Science, 34(4):410–413, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186134.

Anonymous:1967:FMd

[542] Anonymous. Front matter. Philosophy of Science, 34(4):??, December 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186118.

Holton:1968:MPF

[543] Gerald Holton, Edwin C. Kemble, W. V. Quine, S. S. Stevens, and Morton G. White. In memory of Philipp Frank. Philosophy of Science, 35(1): 1–5, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186181.

Hall:1968:BAN

[544] Thomas S. Hall. On biological analogs of Newtonian paradigms. Philosophy of Science, 35(1):6–27, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186182.

Hanna:1968:EE

[545] Joseph F. Hanna. An explication of ’explication’. Philosophy of Science, 35(1):28–44, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186183.

Cornman:1968:MTT

[546] James W. Cornman. Mental terms, theoretical terms, and materialism. Philosophy of Science, 35(1):45–63, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186184.

Korner:1968:BO

[547] S. Körner. On Bergmann’s ontology. Philosophy of Science, 35(1):64– 71, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186185.

Resnik:1968:BRBa

[548] Michael David Resnik. Book review: From Frege to Gödel by Jean van Heijenoort. Philosophy of Science, 35(1):72, March 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186186.

Parsons:1968:BRBa

[549] Howard L. Parsons. Book review: Signification and Significance. A Study of the Relations of Signs and Values by Charles Morris. Philosophy of Science, 35(1):72–73, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186187.

Armour:1968:BRB

[550] Leslie Armour and James R. Horne. Book review: Naturalism and Historical Understanding — Essays on the Philosophy of John Herman Randall, Jr. by John P. Anton. Philosophy of Science, 35(1):73–75, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186188.

Leach:1968:BRBa

[551] James J. Leach. Book review: Foundations of Historical Knowledge by Morton White. Philosophy of Science, 35(1):76–78, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186189.

Anonymous:1968:IA

[552] Anonymous. Important announcement. Philosophy of Science, 35(1): 79, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186190.

Anonymous:1968:AIa

[553] Anonymous. Abstracts from inquiry. Philosophy of Science, 35(1):80– 83, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186191.

Anonymous:1968:ASa

[554] Anonymous. Abstracts from synthese. Philosophy of Science, 35(1):84, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186192.

Anonymous:1968:BMa

[555] Anonymous. Back matter. Philosophy of Science, 35(1):85–99, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186193.

Anonymous:1968:FMa

[556] Anonymous. Front matter. Philosophy of Science, 35(1):??, March 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186180.

Fine:1968:LPQ

[557] Arthur I. Fine. Logic, probability, and quantum theory. Philosophy of Science, 35(2):101–111, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186480.

Humphreys:1968:SAM

[558] W. C. Humphreys. Statistical ambiguity and maximal specificity. Philosophy of Science, 35(2):112–115, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186481.

Hempel:1968:MSL

[559] Carl G. Hempel. Maximal specificity and lawlikeness in probabilistic explanation. Philosophy of Science, 35(2):116–133, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186482.

Rozeboom:1968:NDC

[560] William W. Rozeboom. New dimensions of confirmation theory. Philosophy of Science, 35(2):134–155, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186483.

Webb:1968:MPM

[561] Judson Webb. Metamathematics and the philosophy of mind. Philosophy of Science, 35(2):156–178, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186484.

Berenda:1968:PQS

[562] C. W. Berenda. Phonons — the quantization of sound (and Kant’s second antinomy). Philosophy of Science, 35(2):179–184, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186485.

Marley:1968:AFA

[563] A. A. J. Marley. An alternative “fundamental” axiomatization of multiplicative power relations among three variables. Philosophy of Science, 35(2):185–186, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186486.

Nell:1968:SPS

[564] Edward J. Nell. Strawson, particulars and space. Philosophy of Science, 35(2):187–189, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186487.

Thompson:1968:PSS

[565] John W. Thompson. Polarity in the social sciences and in physics. Philosophy of Science, 35(2):190–194, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186488.

Narveson:1968:BRBa

[566] Jan Narveson. Book review: The Philosophy of Language by Jerrold J. Katz. Philosophy of Science, 35(2):195–197, June 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186489.

Thomas:1968:BRB

[567] Ivo Thomas. Book review: Apuleian Logic, the Nature, Sources, and Influence of Apuleius’s Peri Hermeneias by Mark W. Sullivan. Philosophy of Science, 35(2):197–198, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186490.

Parsons:1968:BRBb

[568] Howard L. Parsons. Book review: The Good City by Lawrence Haworth. Philosophy of Science, 35(2):198, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186491.

Narveson:1968:BRBb

[569] Anne Narveson. Book review: Philosophy, Science and Sense Perception by Maurice Mandelbaum. Philosophy of Science, 35(2):198–199, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186492.

Anonymous:1968:ASb

[570] Anonymous. Abstracts from synthese. Philosophy of Science, 35(2):200– 201, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186493.

Anonymous:1968:ABJ

[571] Anonymous. Abstracts from British journal for the philosophy of science. Philosophy of Science, 35(2):202–203, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186494.

Anonymous:1968:BMb

[572] Anonymous. Back matter. Philosophy of Science, 35(2):??, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186495.

Anonymous:1968:FMb

[573] Anonymous. Front matter. Philosophy of Science, 35(2):??, June 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186479.

Park:1968:QTCa

[574] James L. Park. Quantum theoretical concepts of measurement: Part I. Philosophy of Science, 35(3):205–231, September 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186650.

Hooker:1968:GGH

[575] C. A. Hooker. Goodman, ‘Grue’ and Hempel. Philosophy of Science, 35 (3):232–247, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186651.

See errata [616].

Petrie:1968:SSM

[576] Hugh G. Petrie. The strategy sense of ’methodology’. Philosophy of Science, 35(3):248–257, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186652.

Leach:1968:LS

[577] James J. Leach. The logic of the situation. Philosophy of Science, 35(3): 258–273, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186653.

Hempel:1968:CSC

[578] Carl G. Hempel. On a claim by skyrms concerning lawlikeness and confirmation. Philosophy of Science, 35(3):274–278, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186654.

Coffa:1968:DP

[579] José A. Coffa. Deductive predictions. Philosophy of Science, 35(3):279– 283, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186655.

Hacking:1968:FSS

[580] Ian Hacking. On falling short of strict coherence. Philosophy of Science, 35 (3):284–286, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186656.

Agassi:1968:PTM

[581] Joseph Agassi. Precision in theory and in measurement. Philosophy of Science, 35(3):287–290, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186657.

Lehman:1968:BRB

[582] Hugh Lehman. Book review: Science and Subjectivity by Israel Scheffler. Philosophy of Science, 35(3):291–292, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186658.

Butchvarov:1968:BRB

[583] Panayot Butchvarov. Book review: Experience and Theory. An Essay in the Philosophy of Science by Stephan Körner. Philosophy of Science, 35 (3):292–294, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186659.

Leach:1968:BRBb

[584] James J. Leach. Book review: Skepticism and Historical Knowledge by Jack W. Meiland. Philosophy of Science, 35(3):294–296, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186660.

Maziarz:1968:BRB

[585] Edward A. Maziarz. Book review: The Linguistic Turn: Recent Essays in Philosophical Method by Richard Rorty. Philosophy of Science, 35(3): 296–298, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186661.

vanFraassen:1968:BRB

[586] Bas C. van Fraassen. Book review: The Logical Structure of the World & Pseudo-Problems in Philosophy by Rudolf Carnap; Rolf A. George. Philosophy of Science, 35(3):298–299, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186662.

Wubnig:1968:BRB

[587] Judith Wubnig. Book review: Philosophical Problems in Biology by Vincent E. Smith. Philosophy of Science, 35(3):300–301, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186663.

Anonymous:1968:AIb

[588] Anonymous. Abstracts from inquiry. Philosophy of Science, 35(3):302– 305, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186664.

Anonymous:1968:ASc

[589] Anonymous. Abstracts from synthese. Philosophy of Science, 35(3):306– 307, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186665.

Anonymous:1968:BMc

[590] Anonymous. Back matter. Philosophy of Science, 35(3):??, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186666.

Anonymous:1968:FMc

[591] Anonymous. Front matter. Philosophy of Science, 35(3):??, September 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186649.

Feyerabend:1968:RCC

[592] Paul K. Feyerabend. On a recent critique of complementarity: Part I. Philosophy of Science, 35(4):309–331, December 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186251.

Lacey:1968:CTT

[593] Hugh M. Lacey. The causal theory of time: A critique of Grünbaum’s version. Philosophy of Science, 35(4):332–354, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186252.

Bunge:1968:PTO

[594] Mario Bunge. Physical time: The objective and relational theory. Philosophy of Science, 35(4):355–388, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186253.

Park:1968:QTCb

[595] James L. Park. Quantum theoretical concepts of measurement: Part II. Philosophy of Science, 35(4):389–411, December 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186254.

Byerly:1968:PNC

[596] Henry C. Byerly. Professor Nagel on the cognitive status of scientific theories. Philosophy of Science, 35(4):412–423, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186255.

Resnik:1968:BRBb

[597] Michael David Resnik. Book review: Kleine Schriften by Gottlob Frege; Ignacio Angelelli. Philosophy of Science, 35(4):424–425, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186256.

Bub:1968:BRB

[598] Jeffrey Bub. Book review: Quantum Theory and Reality by Mario Bunge. Philosophy of Science, 35(4):425–429, December 1968. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186257.

Anonymous:1968:VI

[599] Anonymous. Volume information. Philosophy of Science, 35(4):431–434, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186249.

Anonymous:1968:BMd

[600] Anonymous. Back matter. Philosophy of Science, 35(4):??, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186258.

Anonymous:1968:FMd

[601] Anonymous. Front matter. Philosophy of Science, 35(4):??, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186250.

Grunbaum:1969:ICE

[602] Adolf Grünbaum and Wesley C. Salmon. Introduction: The context of these essays. Philosophy of Science, 36(1):1–4, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186559.

Grunbaum:1969:SSC

[603] Adolf Grünbaum. Simultaneity by slow clock transport in the Special Theory of Relativity. Philosophy of Science, 36(1):5–43, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186560.

Salmon:1969:CS

[604] Wesley C. Salmon. The conventionality of simultaneity. Philosophy of Science, 36(1):44–63, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186561.

vanFraassen:1969:CAF

[605] Bas C. van Fraassen. Conventionality in the axiomatic foundations of the Special Theory of Relativity. Philosophy of Science, 36(1):64–73, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186562.

Janis:1969:SST

[606] Allen I. Janis. Synchronism by slow transport of clocks in noninertial frames of reference. Philosophy of Science, 36(1):74–81, March 1969. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186563.

Feyerabend:1969:RCC

[607] Paul K. Feyerabend. On a recent critique of complementarity: Part II. Philosophy of Science, 36(1):82–105, March 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186564.

Caws:1969:BRB

[608] Peter Caws. Book review: Modern Science and Zeno’s Paradoxes by Adolf Grünbaum. Philosophy of Science, 36(1):106–107, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186565.

Anonymous:1969:A

[609] Anonymous. Announcement. Philosophy of Science, 36(1):107–108, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186566.

Anonymous:1969:BMa

[610] Anonymous. Back matter. Philosophy of Science, 36(1):109–126, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186567.

Anonymous:1969:FMa

[611] Anonymous. Front matter. Philosophy of Science, 36(1):??, March 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186558.

Schick:1969:APL

[612] Frederic Schick. Arrow’s proof and the logic of preference. Philosophy of Science, 36(2):127–144, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186166.

Krips:1969:TPQ

[613] H. P. Krips. Two paradoxes in quantum mechanics. Philosophy of Science, 36(2):145–152, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186167.

Corcoran:1969:TLT

[614] John Corcoran. Three logical theories. Philosophy of Science, 36(2):153– 177, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186168.

Kelley:1969:EK

[615] Hillary Jay Kelley. Entropy of knowledge. Philosophy of Science, 36(2): 178–196, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186169.

Anonymous:1969:EGG

[616] Anonymous. Errata: Goodman, ‘Grue’ and Hempel. Philosophy of Science, 36(2):196, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186170.

See [575].

Mayr:1969:FPB

[617] Ernst Mayr. Footnotes on the philosophy of biology. Philosophy of Science, 36(2):197–202, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186171.

Woodson:1969:PEV

[618] M. I. Charles E. Woodson. Parameter estimation vs. hypothesis testing. Philosophy of Science, 36(2):203–204, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186172.

Gordon:1969:SPB

[619] T. J. Gordon and M. J. Raffensperger. A strategy for planning basic research. Philosophy of Science, 36(2):205–218, June 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186173.

Watson:1969:BRB

[620] Richard A. Watson. Book review: Uniformity and Simplicity: A Symposium on the Principle of the Uniformity of Nature by Claude G. Albritton. Philosophy of Science, 36(2):219–221, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186174.

Swanson:1969:BRB

[621] J. W. Swanson. Book review: Conceptual Foundations of Scientific Thought: An Introduction to the Philosophy of Science by Marx Wartofsky. Philosophy of Science, 36(2):221–222, June 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186175.

Michalos:1969:BRBa

[622] Alex C. Michalos. Book review: Public Knowledge by John Ziman. Philosophy of Science, 36(2):222–224, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186176.

Anonymous:1969:ASa

[623] Anonymous. Abstracts from synthese. Philosophy of Science, 36(2):225– 228, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186177.

Anonymous:1969:ABJ

[624] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 36(2):229–232, June 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186178.

Anonymous:1969:BMb

[625] Anonymous. Back matter. Philosophy of Science, 36(2):??, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186179.

Anonymous:1969:FMb

[626] Anonymous. Front matter. Philosophy of Science, 36(2):??, June 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186165.

Hunt:1969:FNM

[627] I. E. Hunt and W. A. Suchting. Force and “natural motion”. Philosophy of Science, 36(3):233–251, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186220.

Causey:1969:DMD

[628] Robert L. Causey. Derived measurement, dimensions, and dimensional analysis. Philosophy of Science, 36(3):252–270, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186221.

Ullian:1969:STT

[629] Joseph S. Ullian. Is any set theory true? Philosophy of Science, 36(3): 271–279, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186222.

Schaffner:1969:CR

[630] Kenneth F. Schaffner. Correspondence rules. Philosophy of Science, 36 (3):280–290, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186223.

Wilson:1969:EANa

[631] Fred Wilson. Explanation in Aristotle, Newton, and Toulmin: Part I. Philosophy of Science, 36(3):291–310, September 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186224.

Williams:1969:GPR

[632] Peter M. Williams. Goodman’s paradox and rules of acceptance. Philosophy of Science, 36(3):311–315, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186225.

Suits:1969:GP

[633] Bernard Suits. Games and paradox. Philosophy of Science, 36(3):316–321, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186226.

Mischel:1969:BRB

[634] Theodore Mischel. Book review: The Philosophy of Science: A Systematic Account by Peter Caws. Philosophy of Science, 36(3):322–324, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186227.

Maziarz:1969:BRB

[635] Edward A. Maziarz. Book review: Problems in the Philosophy of Mathematics by Imre Latakos. Philosophy of Science, 36(3):324–326, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186228.

Michalos:1969:BRBb

[636] Alex C. Michalos. Book review: Philosophy of Science a: Formal Approach by Henry E. Kyburg. Philosophy of Science, 36(3):326–327, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186229.

Gram:1969:BRB

[637] M. S. Gram. Book review: Carnap and Goodman: Two Formalists by Alan Hausman; Fred Wilson. Philosophy of Science, 36(3):327–330, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186230.

Anonymous:1969:BMc

[638] Anonymous. Back matter. Philosophy of Science, 36(3):??, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186231.

Anonymous:1969:FMc

[639] Anonymous. Front matter. Philosophy of Science, 36(3):??, September 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186219.

Massey:1969:TCG

[640] Gerald J. Massey. Toward a clarification of Grünbaum’s conception of an intrinsic metric. Philosophy of Science, 36(4):331–345, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186314.

vanFraassen:1969:MEG

[641] Bas C. van Fraassen. On Massey’s explication of Grünbaum’s conception of metric. Philosophy of Science, 36(4):346–353, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186315.

Evans:1969:FFL

[642] Melbourne G. Evans. On the falsity of the Fitzgerald–Lorentz contraction hypothesis. Philosophy of Science, 36(4):354–362, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186316.

Barrett:1969:CFS

[643] Robert Barrett. On the conclusive falsification of scientific hypotheses. Philosophy of Science, 36(4):363–374, December 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186317.

Wedeking:1969:DQG

[644] Gary Wedeking. Duhem, Quine and Grünbaum on falsification. Philosophy of Science, 36(4):375–380, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186318.

Quinn:1969:SDT

[645] Philip L. Quinn. The status of the D-thesis. Philosophy of Science, 36 (4):381–399, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186319.

Wilson:1969:EANb

[646] Fred Wilson. Explanation in Aristotle, Newton, and Toulmin: Part II. Philosophy of Science, 36(4):400–428, December 1969. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186320.

Grunstra:1969:BRB

[647] Bernard R. Grunstra. Book review: Philosophical Problems of Space and Time by Adolf Grünbaum. Philosophy of Science, 36(4):429–432, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186321.

Anonymous:1969:FMd

[648] Anonymous. Front matter. Philosophy of Science, 36(4):433, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186313.

Anonymous:1969:AD

[649] Anonymous. Abstracts from dialectica. Philosophy of Science, 36(4): 434–435, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186322.

Anonymous:1969:AI

[650] Anonymous. Abstracts from inquiry. Philosophy of Science, 36(4):436– 442, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186323.

Anonymous:1969:ASb

[651] Anonymous. Abstracts from synthese. Philosophy of Science, 36(4):443– 444, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186324.

Anonymous:1969:VI

[652] Anonymous. Volume information. Philosophy of Science, 36(4):445–448, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186312.

Anonymous:1969:BMd

[653] Anonymous. Back matter. Philosophy of Science, 36(4):??, December 1969. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186325.